### v1

# Pruning the Unsurprising: Efficient Code Reasoning via First-Token Surprisal

Abstract
Recently, Large Reasoning Models (LRMs) have demonstrated remarkable capabilities in code reasoning by scaling up the length of Chain-of-Thought (CoT). However, excessively long reasoning traces introduce substantial challenges in terms of training cost, inference latency, and deployment feasibility. While various CoT compression approaches have emerged to address this challenge, they face inherent trade-offs: token-level methods often disrupt syntactic and logical coherence, while step-level methods based on perplexity fail to reliably capture the logically critical reasoning steps. In this paper, we propose ASAP (Anchor-guided, SurprisAl-based Pruning), a novel coarse-to-fine framework for CoT compression. ASAP first performs anchor-guided pruning to preserve the core reasoning structure, which efficiently reduces the search space for subsequent processing. It then enables a logic-aware pruning by selecting logically essential reasoning steps based on a novel first-token surprisal metric. Finally, ASAP teaches models to autonomously generate and leverage these concise CoTs at inference time, enabling efficient reasoning in coding tasks. Experiments show that ASAP achieves state-of-the-art accuracy across multiple code generation benchmarks while substantially reducing training and inference costs. On the challenging LiveCodeBench v4_v5 benchmark, our approach reduces token generation by 23.5% and inference latency by 43.5% compared to the strongest baseline, while achieving a competitive accuracy of 36.19% in Pass@1. Our results highlight a promising direction for building powerful and efficient LRMs. [^1] Code and model available at https://github.com/Zengwh02/ASAP.
Figure 1: Illustration of CoT pruning by ASAP. The original CoT emitted by LRMs contains redundant branches (highlighted in red dashed boxes). ASAP prunes them in the first stage. It then identifies the first tokens of reasoning steps (marked in blue), whose surprisal scores are computed to guide the second-stage pruning.
1 Introduction
Recently, the emergence of Large Reasoning Models (LRMs), including OpenAI’s o1 (Jaech et al. 2024) and DeepSeek-R1 (Guo et al. 2025), has demonstrated remarkable capabilities across a wide range of domains. LRMs empower LLMs with the Chain-of-Thought (CoT) prompting paradigm (Wei et al. 2022), which prompts models to “think step by step”. By generating longer reasoning traces during inference, LRMs are capable of handling complex code reasoning tasks (Shi et al. 2024).
Despite achieving significant performance gains, the current LRMs come at a heavy computational cost when the CoTs scale up. Long CoTs often contain substantial redundancy, particularly pronounced in code generation tasks, where LRMs frequently explore diverse strategies or engage in self-debugging (Bi et al. 2024; Xie et al. 2025; Wu et al. 2025; Qu et al. 2025). This not only increases inference and training costs but also suggests a counterintuitive insight: removing redundant reasoning traces may improve both efficiency and performance.
To improve the efficiency of CoT reasoning, a growing body of research has emerged on efficient reasoning (Qu et al. 2025). A prominent line of work is CoT compression for efficient model fine-tuning, which aims to distill concise yet effective reasoning patterns into the model. For example, TokenSkip (Xia et al. 2025a) adapts general-purpose context compressors such as Selective Context (Li et al. 2023) and LLMLingua-2 (Pan et al. 2024). However, these token-level compression methods are fundamentally ill-suited for code reasoning as they damage the logic of an entire code block within the trace, even if only a variable name is removed. The compressed context makes it hard for models to learn effective reasoning patterns. A more promising direction is step-level pruning. For example, SPIRIT (Cui et al. 2025) prunes the reasoning steps while preserving structural and semantic integrity. However, these approaches face a deeper challenge: How to reliably estimate the logical importance of a reasoning step. SPIRIT relies on perplexity (PPL), but we argue that PPL is an inadequate metric as it primarily reflects linguistic fluency rather than logical necessity and has shown limitations in long-context scenarios (Hu et al. 2024; Fang et al. 2024).
To overcome these challenges, we propose ASAP: A novel framework for Anchor-guided, SurprisAl-based Pruning of CoTs. Our core insight, illustrated in Figure 1, is identifying and pruning redundant steps in CoTs while preserving a coherent logical path. In the first stage, ASAP leverages the inherent structural and logical nature of code to generate a step-by-step reasoning path. The path serves as an anchor for coarse-grained pruning, efficiently identifying and preserving the core reasoning backbone. In the second stage, it enables a fine-grained logic-aware pruning by selecting the first tokens of reasoning steps based on a new First-Token Surprisal metric. The resulting CoT retains only the steps that introduce novel and meaningful information. By fine-tuning on the pruned CoTs, we teach a large reasoning model to generate and leverage these compact reasoning patterns at inference time, enabling efficient reasoning in coding tasks.
We validate our approach through extensive experiments on the DeepSeek-R1-Distill-Qwen-7B and DeepSeek-R1-Distill-Llama-8B (Guo et al. 2025) across a suite of challenging code generation benchmarks, including HumanEval (Chen et al. 2021), HumanEval+ (Liu et al. 2023), LiveCodeBench (Jain et al. 2024), and LeetCodeDataset (Xia et al. 2025b). Our resulting model establishes a new state-of-the-art in the performance-efficiency trade-off. Notably, on the challenging LiveCodeBench v4_v5 benchmark, ASAP achieves 36.19% Pass@1 while reducing token generation by 23.5% and inference latency by 43.5% compared to the strongest baseline.
Our main contributions are summarized as follows:
  • We propose ASAP, a novel two-stage pruning framework that combines coarse-grained anchor-guided pruning with fine-grained surprisal-based pruning to identify and retain logically critical steps.
  • We are the first to propose and empirically validate First-Token Surprisal as a superior metric to perplexity for measuring the logical importance of CoT steps in the code generation task.
  • Extensive experiments on multiple code generation benchmarks demonstrate that models fine-tuned on CoTs pruned by ASAP achieve state-of-the-art accuracy while substantially reducing both training and inference costs.
2 Related Work
2.1 Chain-of-Thought and Advanced Reasoning
Chain-of-Thought (CoT) prompting has emerged as a foundational technique for eliciting reasoning in large language models (LLMs) (Wei et al. 2022). This has sparked a surge of research exploring more sophisticated reasoning frameworks such as Tree of Thoughts (Yao et al. 2023), Graph of Thought (Lei et al. 2023), Natural Program (Ling et al. 2023), etc. More recently, the emergence of powerful Large Reasoning Models (LRMs) such as OpenAI’s o1 (Jaech et al. 2024) and DeepSeek-R1 (Guo et al. 2025) has further advanced the field. These models exhibit strong long-chain reasoning capabilities, which are often enhanced through supervised fine-tuning (SFT) and reinforcement learning (RL) (Jaech et al. 2024; Guo et al. 2025; Kimi et al. 2025; Yang et al. 2025; Yu et al. 2025; Wang et al. 2025), especially in domains such as mathematical and programming reasoning. Unlike prior work that focuses on enhancing reasoning by generating longer or more complex CoTs, we aim to improve both reasoning performance and efficiency by pruning redundant steps from existing CoTs.
Figure 2: The overall framework of ASAP. The process begins with a training sample consisting of a question, an intermediate CoT, and the answer. In Stage 1, an LLM generates a “Direct CoT” ($C_{direct}$) from the question and answer. $C_{direct}$ acts as an anchor to prune the “Original CoT” ($C_{origin}$) into a “Coarse-grained Pruned CoT” ($C_{coarse}$). In Stage 2, we compute the first-token surprisal for each step in $C_{coarse}$ to remove less critical steps, yielding the final “Fine-grained Pruned CoT” ($C^{\prime}$) for fine-tuning.
2.2 Context Compression for LLMs
As tasks become more complex, the context in LLMs has expanded dramatically, resulting in significant computational overhead. This has motivated a line of work on context compression (Li et al. 2025; Fang et al. 2025), aiming to reduce input length while preserving key information. For instance, Selective Context (Li et al. 2023) leverages self-information to delete redundant lexical units. LLMLingua (Jiang et al. 2023) employs an LLM to select key demonstrations and then applies token-level filtering across the context. Its extension, LongLLMLingua (Jiang et al. 2024), introduces document reordering and subsequence reconstruction to enable compression over longer contexts. LLMLingua-2 (Pan et al. 2024) further advances this line by distilling training data and training a token classifier for task-agnostic compression. While these methods have proven effective for text compression, they are not tailored to the structured and logically coherent nature of CoT reasoning. In contrast, our method is explicitly designed for CoT compression, which identifies and preserves logically essential reasoning steps.
2.3 Efficient Reasoning via Fine-Tuning
The increasingly lengthy reasoning traces impose substantial costs, which have motivated a growing body of work focused on efficient reasoning (Qu et al. 2025). Several recent approaches aim to compress CoTs while preserving their effectiveness. C3oT (Kang et al. 2025) uses GPT-4 as a high-quality external compressor to generate shorter CoTs that retain essential reasoning content. TokenSkip (Xia et al. 2025a) enables controllable compression by identifying and removing less important tokens. SPIRIT (Cui et al. 2025) identifies important reasoning steps by detecting perplexity shifts when individual steps are removed. Another paradigm involves compressing CoTs into continuous latent representations, as explored by Coconut (Hao et al. 2024), CCoT (Cheng and Van Durme 2024), CODI (Shen et al. 2025), etc. Our method, ASAP, distinguishes itself from these methods through its explicit two-stage pruning framework, which first generates an anchor for coarse-grained pruning, followed by a fine-grained refinement based on first-token surprisal. By explicitly pruning at the step level, ASAP produces logically condensed human-readable CoTs.
3 Methodology
3.1 Overall Framework
Given a training corpus of triples $\{(Q_{i},\mathbf{C}_{i},A_{i})\}_{i=1}^{N}$—where each $Q_{i}$ denotes a question, $\mathbf{C}_{i}=\{s_{1},\dots,s_{N}\}$ is the intermediate reasoning steps, known as chain-of-thoughts (CoT), generated by a large reasoning model $\mathcal{L}$, and $A_{i}$ represents the predicted answer to the question. Our goal is to compress each $\mathbf{C}_{i}$ into a concise one $\mathbf{C}_{i}^{\prime}$ such that $|\mathbf{C}_{i}^{\prime}|\ll|\mathbf{C}_{i}|$ while the LRM maintains the quality of generated reasoning steps and answers when fine-tuned on the compressed triplets $\{Q_{i},\mathbf{C}_{i}^{\prime},A_{i}\}_{i=1}^{N}$.
We propose ASAP, a novel method for CoT compression, as illustrated in Figure 2. ASAP adopts a coarse-to-fine pruning framework that firstly removes redundant steps from a given reasoning trace and then refines it by selecting the most critical logical steps. In the first stage called Anchor-guided Pruning, ASAP infers a concise, anchor reasoning path called “Direct Thoughts” ($C_{direct}$) based on the $(Q,A)$ pairs. Guided by $C_{direct}$ as an anchor, ASAP prunes the original CoT $C$ into a “coarse-grained pruned CoT” ($C_{coarse}$) using an LLM, which retains the main logical backbone while preliminarily removing digressions and redundancies. In the second stage called Surprisal-based Refining, ASAP performs a logic-aware refinement to $C_{coarse}$. A new metric called First-Token Surprisal is designed to measure the logical importance of a reasoning step and then iteratively filter steps with low surprisal, yielding the final CoT (denoted as $\mathbf{C}^{\prime}$ in our formulation). Finally, all $Q,C^{\prime},A$} tuplets are used to fine-tune a large reasoning model for downstream tasks, which internalizes concise reasoning patterns and substantially improves inference efficiency.
3.2 Anchor-guided Pruning
Generating the Direct CoT Anchor
We prompt an LLM to generate a concise and coherent anchor reasoning trace, $C_{direct}$, based on the question and answer. An anchor is defined as a structured, step-by-step explanation that outlines how to derive the answer from the question. This leverages the inherent logical and procedural nature of code reasoning. The resulting $C_{direct}$ serves as a logical backbone—an anchor—to guide the subsequent pruning of the more elaborate $C_{origin}$. The prompt used for the generation is shown in Appendix E.
Pruning with Pattern Matching
Using $C_{direct}$ as a reference, we prompt the LLM to prune the original CoT $C$. Specifically, the LLM is instructed to: 1) remove unnecessary reasoning steps from $C$; 2) retain all key supporting content that aligns with the logic of $C_{direct}$; and 3) crucially, preserve the original wording without introducing new information. The prompt used for pruning is shown in Appendix E. To mitigate LLM hallucination, we incorporate a validation step based on Gestalt Pattern Matching (Black 2004), which validates structural and semantic alignment with $C$.
Specifically, we design a pattern-matching algorithm that verifies whether each step in $C_{coarse}$ corresponds to a matching step in $C$ while preserving their original order. The matching is performed using Gestalt Pattern Matching (Black 2004) as a text similarity metric. A pruning is considered valid only if all steps in $C_{coarse}$ achieve a similarity score above a predefined threshold $\tau$ when matched against sequential steps in $C$. This guarantees that $C_{coarse}$ forms a consistent substructure of the original reasoning trace. The full pattern-matching algorithm is detailed in Algorithm 1. We leverage high-temperature sampling, which provides the necessary diversity to efficiently re-prompt failed cases, ensuring that a valid $C_{\text{coarse}}$ can be eventually generated.
Algorithm 1 Pattern Matching
1:Original CoT $C$, Coarse-grained Pruned CoT $C_{coarse}$, Threshold $\tau$
2:True if $C_{coarse}$ is valid, False otherwise.
3:function PatternMatch($C,C_{coarse},\tau$)
4:  $S_{origin}\leftarrow\text{SplitStepsByBlankLine}(C)$
5:  $S_{coarse}\leftarrow\text{SplitStepsByBlankLine}(C_{coarse})$
6:  $origin\_idx\leftarrow 0$
7:  for each step $s_{coarse}$ in $S_{coarse}$ do
8:   $found\_match\leftarrow\text{False}$
9:   while $origin\_idx<\text{Length}(S_{origin})$ do
10:     $s_{origin}\leftarrow S_{origin}[origin\_idx]$
11:     $score\leftarrow\text{GestaltSimilarity}(s_{origin},s_{coarse})$
12:     if $score\geq\tau$ then
13:      $found\_match\leftarrow\text{True}$
14:      $origin\_idx\leftarrow origin\_idx+1$
15:      break
16:     end if
17:     $origin\_idx\leftarrow origin\_idx+1$
18:   end while
19:   if not found_matchfound\_match then
20:     return False
21:   end if
22:  end for
23:  return True
24:end function
3.3 Surprisal-based Refining
Following the coarse-grained pruning, we perform a meticulous, logic-aware refinement in $C_{coarse}$ to identify more subtle redundancies within the core reasoning path.
First-Token Surprisal as Logical Importance
We introduce First-Token Surprisal as a novel metric to precisely quantify the logical importance of each step, enabling us to iteratively filter out the least informative ones and produce the final highly condensed CoT.
In information theory, entropy is used to quantify the average level of uncertainty or information associated with a variable’s potential states or possible outcomes (Shannon 1948). In the era of LLMs, this concept can be applied to each token generated by the model. Given an LLM generates a sequence of tokens $x_{1},x_{2},\dots,x_{N}$, the conditional probability of the $n$-th token is given by $p(x_{n}|x_{<n})$, derived from the model’s logits. Token-level entropy measures the model’s uncertainty over the entire vocabulary $V$ for the next token (Malinin and Gales 2020; Kuhn, Gal, and Farquhar 2023). It is the surprisal over all possible outcomes:
$H(x_{n}|x_{<n})=-\sum_{v\in V}p(v|x_{<n})\log p(v|x_{<n})$ (1)
In large reasoning models, the first token of each step often signals a logical transition or continuation. High entropy indicates the model’s uncertainty, suggesting that the step may introduce semantically or logically significant information. Conversely, low entropy typically reflects predictable continuations or syntactic fillers, implying limited contribution to the overall reasoning process (Wang et al. 2025; Cheng et al. 2025).
However, since the actual next token is known in the sequence, we directly compute the surprisal of the target token, a more efficient and targeted metric that quantifies the unexpectedness of the target token $x_{t}$:
$S(x_{t}|x_{<n})=-\log p(x_{t}|x_{<n})$ (2)
Surprisal quantifies how unexpected a given token is to the model. We focus specifically on the first-token surprisal of each reasoning step. A high surprisal value at the beginning of a step indicates logical salience—that is, the step either initiates a new line of reasoning or introduces novel information. In contrast, a low surprisal value suggests a continuation of prior context or redundant elaboration. Accordingly, we prune low-surprisal steps to obtain more concise and informative reasoning traces.
Iterative Pruning Algorithm
Based on the first-token surprisal metric, we propose a length-controllable iterative pruning algorithm. As shown in Algorithm 2, it calculates the surprisal score for the first token of each reasoning step within the $C_{coarse}$. It then iteratively removes the step with the lowest surprisal score until the total length of the CoT falls below a predefined token budget $L_{max}$. This process yields the final $C^{\prime}$, which is highly compressed while retaining the most logically significant steps.
Algorithm 2 Iterative Pruning via First-Token Surprisal
1:Coarse-grained Pruned CoT $C_{coarse}$, Max Tokens $L_{max}$, Model $M$, Tokenizer $T$
2:Fine-grained Pruned CoT $C^{\prime}$
3:function FineGrainedPrune($C_{coarse},L_{max},M,T$)
4:  if $\text{Length}(T(C_{coarse}))\leq L_{max}$ then
5:   return $C_{coarse}$
6:  end if
7:  $S\leftarrow\text{SplitStepsByBlankLine}(C_{coarse})$
8:  $SurprisalScores\leftarrow\text{CalculateAll}(S,M,T)$
9:  $StepsToPrune\leftarrow\text{SortByScore}(S,\text{SurprisalScores})$
10:  $S_{current}\leftarrow S$
11:  for each step $s_{prune}$ in StepsToPruneStepsToPrune do
12:   $S_{temp}\leftarrow S_{current}\setminus\{s_{prune}\}$
13:   $C_{temp}\leftarrow\text{Join}(S_{temp})$
14:   if $\text{Length}(T(C_{temp}))\leq L_{max}$ then
15:     $S_{current}\leftarrow S_{temp}$
16:     break
17:   end if
18:   $S_{current}\leftarrow S_{temp}$
19:  end for
20:  $C^{\prime}\leftarrow\text{Join}(S_{current})$
21:  return $C^{\prime}$
22:end function
3.4 Supervised Fine-tuning
Following the two-stage pruning by our ASAP framework, we construct the final training dataset using the compressed triplets $\{(Q_{i},\mathbf{C}_{i}^{\prime},A_{i})\}_{i=1}^{N}$. For each instance, we concatenate the pruned CoT ($\mathbf{C}_{i}^{\prime}$) and the final answer ($A_{i}$) to form the complete target response, denoted as $R_{i}$. We then fine-tune a large reasoning model on this dataset. The training goal is to minimize the negative log-likelihood of the target response tokens, conditioned on the input question. Formally, the loss is defined as:
$\mathcal{L}=-\sum_{i=1}^{N}\sum_{j=1}^{|R_{i}|}\log P_{\theta}(r_{i,j}|Q_{i},r_{i,<j})$ (3)
where $r_{i,j}$ is the $j$-th token of the target response $R_{i}$, and $\theta$ represents the parameters of the model being fine-tuned. This supervised fine-tuning process effectively distills the knowledge from our pruning framework into the model. By training on these compact and logically salient examples, the model learns to internalize the efficient reasoning patterns.
4 Experiments
Methods HumanEval LiveCodeBench LeetCodeDataset
HE HE+ v1_v3 v4_v5 All
Acc $\uparrow$ Tok $\downarrow$ Lat $\downarrow$ Acc $\uparrow$ Tok $\downarrow$ Lat $\downarrow$ Acc $\uparrow$ Tok $\downarrow$ Lat $\downarrow$ Acc $\uparrow$ Tok $\downarrow$ Lat $\downarrow$ Acc $\uparrow$ Tok $\downarrow$ Lat $\downarrow$
Zero-shot 73.78 3051 1.16 68.29 3051 1.16 42.16 7088 3.59 25.37 8336 5.15 19.74 8680 4.95
Original 78.05 2973 1.12 75.61 2973 1.12 52.12 6611 3.15 30.97 8289 4.83 25.00 8485 4.72
Selective Context 59.76 2979 1.13 54.88 2979 1.13 30.23 7025 3.75 16.79 8558 5.35 15.79 8461 4.90
LLMLingua-2 71.34 3075 1.19 68.29 3075 1.19 38.89 6953 3.60 22.76 8474 5.31 17.54 8513 4.81
TokenSkip 76.83 2823 1.07 73.78 2823 1.07 32.35 7095 3.85 20.15 8400 5.37 18.42 8503 4.87
SPIRIT 83.54 2764 1.07 75.61 2764 1.07 50.82 6524 3.09 33.58 7892 4.62 25.00 8186 4.45
ASAP 84.15 2464 0.98 78.66 2464 0.98 54.74 5177 2.09 36.19 6035 2.61 27.63 7541 3.48
Table 1: Experimental results of different methods with DeepSeek-R1-Distill-Qwen-7B. We report accuracy (Acc), average number of generated tokens (Tok), and average generation latency (Lat) measured in seconds. The best results are highlighted in bold, and the second-best are underlined.
4.1 Experimental Setup
Models and Datasets
All experiments are conducted on the DeepSeek-R1-Distill-Qwen-7B and DeepSeek-R1-Distill-Llama-8B (Guo et al. 2025), with DeepSeek-R1-Distill-Qwen-7B as the default backbone across all settings. We use the Python subset of the CodeForces-CoTs (Hugging Face 2025) dataset for training. The dataset consists of high-quality Chain-of-Thought (CoT) samples generated by DeepSeek-R1, making it particularly suitable as training data for competitive programming reasoning tasks.
Benchmarks
We evaluate our method on a suite of widely used code generation benchmarks that span a range of difficulty levels:
  • HumanEval (Chen et al. 2021) consists of 164 hand-crafted programming problems. HumanEval+ (Liu et al. 2023) extends HumanEval with approximately 80 times more unique test cases.
  • LiveCodeBench (Jain et al. 2024) is a comprehensive, contamination-free benchmark that continuously aggregates new problems from competitive programming platforms. We use two subsets: v1_v3 (612 problems) and v4_v5 (268 problems).
  • LeetCodeDataset (Xia et al. 2025b) provides a high-quality evaluation set of 228 problems collected from LeetCode. It covers over 90% of Python problems on the platform, with each problem featuring over 100 test cases.
Baselines
We compare our method against a comprehensive set of baselines:
  • Zero-shot: The original model without any task-specific fine-tuning.
  • Original: The model is fine-tuned on the original, uncompressed CoTs from the training data.
  • Selective Context (Li et al. 2023): Identifies and prunes redundant lexical units in the input using self-information, aiming to retain only the most informative tokens.
  • LLMLingua-2 (Pan et al. 2024): Distills GPT-4’s token importance signals into a lightweight Transformer encoder, which is trained as a token classifier. This approach compresses content using the token classifier.
  • TokenSkip (Xia et al. 2025a): Trains a model to selectively skip less important tokens in a CoT, allowing for controllable compression based on token importance scores.
  • SPIRIT (Cui et al. 2025): Identifies critical reasoning steps by measuring perplexity shifts—steps are retained if their removal significantly increases the perplexity of the remaining text.
Except for the zero-shot baseline, all other methods involve fine-tuning the model using CoTs that have been processed according to their strategies.
Metrics
To comprehensively evaluate each method, we report both accuracy and inference efficiency:
  • Accuracy is measured by Pass@1, which denotes the percentage of problems correctly solved on the first attempt.
  • Efficiency is evaluated using two metrics: the average number of generated tokens and the average generation latency measured in seconds per sample.
Implementation Details
All models are fully fine-tuned (i.e., full-parameter fine-tuning) using Unsloth (Daniel Han and team 2023) for efficiency. For inference, we utilize the vLLM (Kwon et al. 2023) engine to enable fast and efficient decoding. To ensure reproducibility, we use greedy decoding and disable prefix caching. The default generation token budget is adjusted based on task difficulty: 6K for HumanEval and HumanEval+, and 10K for LiveCodeBench and LeetCodeDataset. All experiments are conducted on NVIDIA H20 GPUs. Additional implementation details are provided in Appendix A.
4.2 Main Results
The comprehensive results of our method compared to all baselines are presented in Table 1. The results show that the model fine-tuned on CoTs pruned by ASAP consistently achieves the best trade-off between accuracy and efficiency. It outperforms all baselines in terms of accuracy, while generating the fewest tokens and achieving the lowest generation latency across all benchmarks. Statistical analysis using the Wilcoxon signed-rank test confirms that the improvements of our method over all baselines are statistically significant ($p<0.05$).
A deeper analysis of the results reveals a clear distinction between token-level and step-level pruning strategies. Token-level baselines like Selective Context, LLMLingua-2, and TokenSkip exhibit a significant performance degradation compared to the original CoTs. This is because the token removal disrupts the syntactic structure and semantic coherence of the original reasoning steps. Consequently, the fine-tuning data becomes fragmented and grammatically unnatural, making it difficult for the model to learn the intended logical flow of the CoT. Step-level methods, such as SPIRIT, perform significantly better than token-level pruning methods, due to the preservation of sentence-level integrity. While SPIRIT improves efficiency over the Original with comparable accuracy, our method achieves higher efficiency and accuracy at the same time. This improvement is particularly pronounced on the challenging LiveCodeBench v4_v5 benchmark: ASAP reduces the average number of generated tokens by 23.5% (from 7892 to 6035) and lowers generation latency by 43.5% (from 4.62s to 2.61s), while also achieving a 7.8% improvement in accuracy (Pass@1 increases from 33.58% to 36.19%). These results verify our core hypothesis that pruning reasoning chains based on first-token surprisal is more effective than perplexity or token-level pruning.
4.3 Ablation Study
To validate the contribution and necessity of each component in our two-stage pruning framework, we conduct a detailed ablation study. Specifically, we evaluate the following three variants:
  • ASAP w/o Coarse-grained Pruning: This variant skips Stage 1 and applies only the fine-grained surprisal-based pruning directly to the original CoT.
  • ASAP w/o Fine-grained Pruning: This variant applies only the coarse-grained anchor-guided pruning from Stage 1, omitting the surprisal-based refinement in Stage 2.
  • ASAP w/o Any Pruning: Equivalent to the Original baseline, where the model is fine-tuned on the full, uncompressed CoT without any pruning.
Table 2 reports results on LiveCodeBench v4_v5, which is representative of the consistent trends observed across benchmarks. Additional results are included in Appendix B.
The ablation study demonstrates that both pruning stages are essential and mutually complementary for achieving optimal accuracy and efficiency. First, removing the coarse-grained pruning stage (w/o Coarse-grained Pruning) leads to a drop in both accuracy and efficiency. While the accuracy decrease is modest, the generation latency increases by a substantial 76.2% (from 2.61s to 4.60s), underscoring the importance of stage 1 in eliminating redundancy and narrowing the search space for subsequent processing. Second, removing the fine-grained pruning stage (w/o Fine-grained Pruning) results in a significant degradation across all metrics. The accuracy drops by 12.4% (Pass@1 decreases from 36.19% to 31.72%) relative to the ASAP, and efficiency improvements are largely lost. This highlights that Stage 2—our surprisal-based pruning mechanism—is central to precisely identifying and preserving the most critical logical steps. In summary, the two stages operate in synergy: the coarse-grained anchor-guided pruning preliminarily removes redundant steps, while the fine-grained surprisal-based pruning further extracts the logical core of the reasoning trace. This two-stage process is key to ASAP’s strong performance, significantly outperforming any single-stage variant.
Variants Acc $\uparrow$ Tok $\downarrow$ Lat $\downarrow$
ASAP 36.19 6035 2.61
w/o Coarse-grained Pruning 35.07 7735 4.60
w/o Fine-grained Pruning 31.72 8061 4.83
w/o Any Pruning 30.97 8289 4.83
Table 2: Ablation study of different pruning strategies for ASAP on LiveCodeBench v4_v5. We report accuracy (Acc), average number of generated tokens (Tok), and average generation latency (Lat) measured in seconds.
4.4 Impact of Token Budget
To evaluate the performance scalability and resource sensitivity of our method, we analyze its behavior under varying inference-time token budgets (i.e., the maximum number of tokens the model is allowed to generate). We compare ASAP against the three strong baselines—SPIRIT, Original, and Zero-shot—across all benchmarks, and observe consistent trends. For clarity, we present results on LiveCodeBench v4_v5 under six budget settings ranging from 2K to 12K tokens. Results for other benchmarks and additional efficiency statistics are provided in Appendix C.
As shown in Figure 3, our findings highlight three key advantages of ASAP: First, ASAP consistently outperforms all baselines across all budget settings on the benchmark. This demonstrates the robustness of our method under different resource constraints. Second, ASAP exhibits smooth performance scaling with respect to the token budget. Unlike baseline methods such as Zero-shot, whose accuracy can fluctuate as the budget increases (e.g., Zero-shot shows a decline from 8K to 10K), ASAP reliably improves with more available tokens. This predictable scaling behavior enables users to confidently configure budget limits to meet specific accuracy or latency targets. Third, ASAP achieves superior performance-efficiency trade-offs. For example, ASAP with just an 8K token budget achieves higher accuracy than SPIRIT and Original at a much larger 12K budget. These results further validate the practical utility of ASAP in real-world scenarios with constrained resources and varying performance requirements.
Figure 3: Performance of ASAP on LiveCodeBench v4_v5 under different token budgets.
4.5 Training Efficiency
A central motivation for CoT compression is to reduce the computational burden during training and inference, specifically shortening the lengthy reasoning traces typically required by large reasoning models. Fine-tuning on compressed CoTs not only lowers training-time resource consumption but also helps the model internalize efficient reasoning patterns. To this end, we aim to achieve a high compression ratio across all methods while preserving the core logical steps. More details are provided in Appendix A.
To quantify the training efficiency gains, we report two key metrics in Table 3: the average number of tokens per sample and the average training time measured in seconds per step. The results highlight the training efficiency advantage of ASAP. By generating the most compact yet logically rich CoTs, our approach significantly reduces training overhead. Compared to the uncompressed baseline (Original), our method reduces the number of training tokens by 75.6% and shortens training time by 60.7%. These savings substantially exceed those achieved by all other baselines. ASAP not only achieves a more favorable accuracy-efficiency trade-off at inference time, but also enables a more resource-efficient training process, making it a practical and cost-effective solution for real-world deployment.
Methods Tokens Time
Original 13023 80.11
Selective Context 6722 (-48.4%) 63.41 (-20.9%)
LLMLingua-2 6919 (-46.9%) 65.25 (-18.6%)
TokenSkip 9813 (-24.6%) 77.27 (-3.6%)
SPIRIT 6082 (-53.3%) 57.45 (-28.3%)
ASAP 3178 (-75.6%) 31.48 (-60.7%)
Table 3: Training efficiency comparison. We report the average number of tokens per sample and training time measured in seconds per step. Percentages indicate the reduction relative to the Original baseline.
4.6 Generalization to Different Architectures
To validate the generalizability of ASAP, we replicate our main experiments on the DeepSeek-R1-Distill-Llama-8B. Following the same experimental protocol, we compare ASAP against three strong baselines: Zero-shot, Original, and SPIRIT. We observe consistent trends across all benchmarks, so for brevity, we present representative results on two key benchmarks: LiveCodeBench v4_v5 and LeetCodeDataset in Table 4. The results across all benchmarks are reported in the Appendix D.
The results on DeepSeek-R1-Distill-Llama-8B are highly consistent with our findings on the DeepSeek-R1-Distill-Qwen-7B, confirming that the effectiveness of ASAP generalizes beyond a specific model family. As shown in the Table 4, ASAP achieves the highest accuracy on both benchmarks, and the improvements in efficiency are even more pronounced. On LiveCodeBench, for instance, ASAP not only surpasses the accuracy of the Original baseline (32.84% vs. 31.34%) but also generates 49.1% fewer tokens and reduces latency by over 3x (from 8.60s to 2.69s). Compared to the Zero-shot baseline, our method reduces token generation by 50.9% and latency by a remarkable 69.8%. This suggests that ASAP is particularly effective in identifying and distilling the core reasoning patterns, validating its robustness and broad applicability for improving reasoning efficiency across different model families.
Methods LiveCodeBench LeetCodeDataset
Acc $\uparrow$ Tok $\downarrow$ Lat $\downarrow$ Acc $\uparrow$ Tok $\downarrow$ Lat $\downarrow$
Zero-shot 25.00 8508 8.90 27.19 8358 8.65
Original 31.34 8202 8.60 26.32 8413 8.85
SPIRIT 30.22 7913 8.45 26.75 8449 8.73
ASAP 32.84 4175 2.69 27.63 3792 2.42
Table 4: Experimental results of different methods with DeepSeek-R1-Distill-Llama-8B. We report accuracy (Acc), average number of generated tokens (Tok), and average generation latency (Lat) measured in seconds. The best results are highlighted in bold.
5 Conclusion
In this paper, we addressed the challenge of balancing performance and efficiency in Chain-of-Thought (CoT) reasoning, where longer reasoning traces improve accuracy but introduce significant redundancy and computational overhead. We proposed ASAP, a novel two-stage pruning framework that first generates an anchor for coarse-grained pruning, followed by a fine-grained refinement based on first-token surprisal. Extensive experiments demonstrate that models fine-tuned on CoTs pruned by ASAP achieve state-of-the-art results on multiple code generation benchmarks while substantially reducing both training and inference costs. Although our current work focuses on competitive programming tasks within the code generation domain, we believe the framework is general and broadly applicable.
Appendix A Implementation Details
Software and Hardware
For fine-tuning, we utilized the unsloth library [^2] https://pypi.org/project/unsloth/2025.5.6/ for its memory-efficient optimizations. For inference, we employed the vLLM engine [^3] https://pypi.org/project/vllm/0.8.4/ to maximize throughput and efficiency. All experiments were conducted on NVIDIA H20 GPUs and Intel Xeon Platinum 8480+ CPUs.
Fine-tuning Configuration
We performed full-parameter fine-tuning for all models in our experiments. Key hyperparameters included precision set to bf16, num_train_epochs set to 10, and a max_seq_length of 16384. We used a per_device_train_batch_size of 1 with gradient_accumulation_steps set to 16, resulting in an effective batch size of 16. For the optimizer, we used AdamW with a cosine_with_min_lr learning rate scheduler. The warmup_ratio was set to 0.03, and the scheduler’s min_lr_rate was 0.1 of the peak learning rate. To stabilize training, we applied gradient clipping with a max_grad_norm of 0.2. Based on preliminary experiments, we set the peak learning rate to $4\times 10^{-5}$ for the DeepSeek-R1-Distill-Qwen-7B and $2\times 10^{-5}$ for the DeepSeek-R1-Distill-Llama-8B. Due to the high computational cost of full-parameter fine-tuning, the model is fine-tuned by a single run with a fixed random seed 42.
Inference and Evaluation Protocol
All inference benchmarks were run using the vLLM engine with dtype set to bfloat16 and gpu_memory_utilization set to 0.9. To ensure deterministic and reproducible outputs, we set the sampling temperature to 0.0 and set enable_prefix_caching to False.
Baseline Details
Following established practices, we used a consistent scoring model; as our primary model is DeepSeek-R1-Distill checkpoints, we employed DeepSeek-R1-Distill-Qwen-7B for all model-scoring tasks. To ensure a fair comparison, we standardize the input format across all methods by preserving the original question and final answer, and applying compression only to the CoT reasoning steps. To balance compression ratio and content retention, we set the target compression ratio to 0.5 for all baseline methods, except for TokenSkip, where we follow its original design that allows a controllable compression ratio between 0.5 and 1.0. Additionally, since the original SPIRIT method is computationally expensive when applied to extremely long CoTs, we adopt a modified version to ensure fair comparison: specifically, we compute perplexity once per reasoning step and iteratively remove steps until the target ratio is met. This variant retains the core idea of SPIRIT while improving scalability in our evaluation setting.
Hyperparameters for Our Method
Our method involves several stages. For the LLM-guided Coarse-grained Pruning stage, we employed DeepSeek-V3 for economic reasons. When generating the Direct CoT, we used a deterministic setting (temperature=0.0, top_p=1.0), while for making the final pruning result, we increased exploration (temperature=2.0, top_p=1.0). For Pattern Matching, the similarity threshold $\tau$ was set to 0.6. Finally, during Surprisal-based Fine-grained Pruning, the maximum token budget was set to 4096 to ensure a deep level of compression.
Appendix B Ablation Study of Different Pruning Strategies
To validate the contribution and necessity of each component in our two-stage pruning framework, we conduct a detailed ablation study. Specifically, we evaluate the following three variants: ASAP w/o Coarse-grained Pruning, ASAP w/o Fine-grained Pruning and ASAP w/o Any Pruning. We present results on the HumanEval, HumanEval+, LiveCodeBench v1_v3, and LeetCodeDatsets benchmarks in Table 5, Table 6, Table 7, and Table 8.
Variants Acc $\uparrow$ Tok $\downarrow$ Lat $\downarrow$
ASAP 84.15 2464 0.98
w/o Coarse-grained Pruning 82.32 2839 1.10
w/o Fine-grained Pruning 71.34 2897 1.10
w/o Any Pruning 78.05 2973 1.12
Table 5: Ablation study of different pruning strategies for ASAP on HumanEval. We report accuracy (Acc), average number of generated tokens (Tok), and average generation latency (Lat) measured in seconds.
Variants Acc $\uparrow$ Tok $\downarrow$ Lat $\downarrow$
ASAP 78.66 2464 0.98
w/o Coarse-grained Pruning 78.05 2839 1.10
w/o Fine-grained Pruning 67.07 2897 1.10
w/o Any Pruning 75.61 2973 1.12
Table 6: Ablation study of different pruning strategies for ASAP on HumanEval+. We report accuracy (Acc), average number of generated tokens (Tok), and average generation latency (Lat) measured in seconds.
Variants Acc $\uparrow$ Tok $\downarrow$ Lat $\downarrow$
ASAP 54.74 5177 2.09
w/o Coarse-grained Pruning 53.92 6107 2.77
w/o Fine-grained Pruning 51.14 6599 3.20
w/o Any Pruning 52.12 6611 3.15
Table 7: Ablation study of different pruning strategies for ASAP on LiveCodeBench v1_v3. We report accuracy (Acc), average number of generated tokens (Tok), and average generation latency (Lat) measured in seconds.
Variants Acc $\uparrow$ Tok $\downarrow$ Lat $\downarrow$
ASAP 27.63 7541 3.48
w/o Coarse-grained Pruning 24.12 7954 3.75
w/o Fine-grained Pruning 25.44 8326 4.77
w/o Any Pruning 25.00 8485 4.72
Table 8: Ablation study of different pruning strategies for ASAP on LeetCodeDataset. We report accuracy (Acc), average number of generated tokens (Tok), and average generation latency (Lat) measured in seconds.
Appendix C Performance under Different Token Budgets
To evaluate the performance scalability and resource sensitivity of our method, we analyze its behavior under varying inference-time token budgets (i.e., the maximum number of tokens the model is allowed to generate). We compare ASAP with three strong baselines—SPIRIT, Original, and Zero-shot—on HumanEval, LiveCodeBench, and LeetcodeDataset. For HumanEval, we evaluate the performance under four budget settings, ranging from 1K to 6K tokens. For LiveCodeBench and LeetCodeDataset, we evaluate the performance under six budget settings, ranging from 2K to 12K tokens. Results are shown in Table 9, Table 10 and Table 11.
Budget Zero-Shot Original SPIRIT ASAP
Acc Tok Lat Acc Tok Lat Acc Tok Lat Acc Tok Lat
1K 10.37 1007 0.28 15.85 983 0.28 11.59 995 0.28 25.00 946 0.27
2K 45.12 1813 0.53 44.51 1702 0.49 49.39 1690 0.49 56.71 1502 0.44
4K 71.95 2561 0.85 67.68 2511 0.82 76.83 2401 0.80 75.00 2116 0.72
6K 73.78 3051 1.16 78.05 2973 1.12 83.54 2764 1.07 84.15 2464 0.98
(a) Results on HumanEval
Budget Zero-Shot Original SPIRIT ASAP
Acc Tok Lat Acc Tok Lat Acc Tok Lat Acc Tok Lat
1K 9.76 1007 0.28 14.63 983 0.28 10.98 995 0.28 23.78 946 0.27
2K 42.68 1813 0.53 43.29 1702 0.49 47.56 1690 0.49 54.88 1502 0.44
4K 66.46 2561 0.85 65.85 2511 0.82 69.51 2401 0.80 71.34 2116 0.72
6K 68.29 3051 1.16 75.61 2973 1.12 75.61 2764 1.07 78.66 2464 0.98
(b) Results on HumanEval+
Table 9: Results of different methods under different budgets on HumanEval. We report accuracy (Acc), average number of generated tokens (Tok), and average generation latency (Lat) measured in seconds.
. Budget Zero-shot Original SPIRIT ASAP Acc Tok Lat Acc Tok Lat Acc Tok Lat Acc Tok Lat 2K 16.50 1966 0.52 17.16 1920 0.51 18.95 1908 0.51 21.57 1833 0.49 4K 32.68 3499 1.06 30.72 3432 1.05 34.80 3370 1.03 34.97 3244 1.00 6K 39.05 4806 1.70 42.65 4673 1.67 43.14 4605 1.64 46.24 4358 1.54 8K 44.28 5903 2.46 47.71 5723 2.43 51.80 5515 2.27 52.61 4919 1.90 10K 42.16 7088 3.59 52.12 6611 3.15 50.82 6524 3.09 54.74 5177 2.09 12K 43.95 7988 5.10 54.41 7473 4.22 51.63 7362 4.09 55.56 5322 2.27
(a) Results on LiveCodeBench v1_v3
Budget Zero-shot Original SPIRIT ASAP
Acc Tok Lat Acc Tok Lat Acc Tok Lat Acc Tok Lat
2K 6.72 2021 0.59 6.34 1999 0.57 8.21 1993 0.56 13.43 1930 0.54
4K 16.79 3820 1.22 15.67 3799 1.20 20.15 3712 1.18 20.90 3594 1.15
6K 23.13 5444 2.07 22.76 5397 2.00 26.49 5237 1.93 30.60 4988 1.85
8K 25.37 6927 3.27 25.74 6882 3.24 30.60 6634 3.09 35.07 5793 2.38
10K 25.37 8336 5.15 30.97 8289 4.83 33.58 7892 4.62 36.19 6035 2.61
12K 25.75 9706 7.44 32.46 9567 7.10 34.33 8987 6.73 36.57 6128 2.76
(b) Results on LiveCodeBench v4_v5
Table 10: Results of different methods under different budgets on LiveCodeBench. We report accuracy (Acc), average number of generated tokens (Tok), and average generation latency (Lat) measured in seconds.
Budget Zero-Shot Original SPIRIT ASAP
Acc Tok Lat Acc Tok Lat Acc Tok Lat Acc Tok Lat
2K 7.02 2028 0.53 6.14 2020 0.53 7.02 2001 0.53 10.09 1965 0.53
4K 13.16 3848 1.21 13.16 3854 1.21 16.23 3789 1.19 15.79 3758 1.19
6K 16.23 5553 2.04 16.67 5548 2.04 18.86 5407 2.00 19.30 5387 2.00
8K 19.30 7165 3.27 22.37 7104 3.18 22.37 6882 3.04 23.25 6722 2.88
10K 19.74 8680 4.95 25.00 8485 4.72 25.00 8186 4.45 27.63 7541 3.48
12K 21.49 10142 7.58 28.07 9717 7.09 26.32 9354 6.86 27.63 7902 3.83
Table 11: Results of different methods under different budgets on LeetCodeDataset. We report accuracy (Acc), average number of generated tokens (Tok), and average generation latency (Lat) measured in seconds.
Appendix D Generalization to Different Architectures
To evaluate the generalizability of ASAP, we replicate our main experiments on the DeepSeek-R1-Distill-Llama-8B. Following the same experimental protocol, we compare ASAP against three baselines: Zero-shot, Original, and SPIRIT. Results on the HumanEval, HumanEval+, LiveCodeBench v1_v3, LiveCodeBench v4_v5, and LeetCodeDataset benchmarks are shown in Table 12.
Methods HumanEval LiveCodeBench LeetCodeDataset
HE HE+ v1_v3 v4_v5 All
Acc $\uparrow$ Tok $\downarrow$ Lat $\downarrow$ Acc $\uparrow$ Tok $\downarrow$ Lat $\downarrow$ Acc $\uparrow$ Tok $\downarrow$ Lat $\downarrow$ Acc $\uparrow$ Tok $\downarrow$ Lat $\downarrow$ Acc $\uparrow$ Tok $\downarrow$ Lat $\downarrow$
Zero-shot 68.29 3334 1.86 64.02 3334 1.86 44.12 7162 6.92 25.00 8508 8.90 27.19 8358 8.65
Original 82.32 2978 1.63 76.22 2978 1.63 52.61 6614 6.16 31.34 8202 8.60 26.32 8413 8.85
SPIRIT 76.22 3159 1.74 72.56 3159 1.74 52.61 6280 5.84 30.22 7913 8.45 26.75 8449 8.73
ASAP 80.49 2494 1.30 76.83 2494 1.30 48.86 3605 2.18 32.84 4175 2.69 27.63 3792 2.42
Table 12: Experimental results of different methods with DeepSeek-R1-Distill-Llama-8B. We report accuracy (Acc), average number of generated tokens (Tok), and average generation latency (Lat) measured in seconds.
Appendix E Prompts Used in Our Method
Prompt for Direct CoT Generation Given a question, please tell me how to get this answer step by step. Question: ``` {question} ``` Answer: ```python {answer} ``` Only return a detailed (with code) step-by-step solution (containing only “Step-by-Step Solution” and “Final Code”). The detailed step-by-step solution is:
Prompt for Coarse-grained Pruning Compress the given thinking by referring to the provided solution. The goal is to remove irrelevant reasoning paths while retaining all content along the core reasoning path. Compression must be based on thinking, ensuring that the original wording and structure are preserved as much as possible. Follow these strict rules: 1. Use thinking as the foundation: Do not rewrite or replace its content with solution——only use solution to determine which parts are relevant. 2. Remove unnecessary reasoning: Aggressively remove alternative paths that are not part of the core reasoning path. 3. Retain key supporting content: Keep examples, reflections, and tests that help illustrate, verify, or analyze the core reasoning path. 4. Preserve original words: Do not paraphrase, reorder, or change any words. 5. Do not add new words: Do not introduce new concepts, symbols, or abbreviations. If you understand, compress the following thinking based on the given solution. Solution: ``` {solution} ``` Thinking: ``` {think} ``` The compressed thinking is:

### v2

@@
-Pruning the Unsurprising: Efficient Code Reasoning via First-Token Surprisal
+Pruning the Unsurprising: Efficient LLM Reasoning via First-Token Surprisal
 Abstract
-Recently, Large Reasoning Models (LRMs) have demonstrated remarkable capabilities in code reasoning by scaling up the length of Chain-of-Thought (CoT).
-However, excessively long reasoning traces introduce substantial challenges in terms of training cost, inference latency, and deployment feasibility.
-While various CoT compression approaches have emerged to address this challenge, they face inherent trade-offs: token-level methods often disrupt syntactic and logical coherence, while step-level methods based on perplexity fail to reliably capture the logically critical reasoning steps.
+Large Reasoning Models (LRMs) have demonstrated remarkable capabilities by scaling up the length of Chain-of-Thought (CoT).
+However, excessively long reasoning traces pose substantial challenges for training cost and inference latency.
+While various CoT compression approaches have emerged to address this challenge, they face inherent trade-offs: token-level methods often disrupt syntactic and logical coherence, while step-level methods based on perplexity fail to reliably capture the logically critical reasoning steps because of the dilution of logical information.
 In this paper, we propose ASAP (Anchor-guided, SurprisAl-based Pruning), a novel coarse-to-fine framework for CoT compression.
 ASAP first performs anchor-guided pruning to preserve the core reasoning structure, which efficiently reduces the search space for subsequent processing.
-It then enables a logic-aware pruning by selecting logically essential reasoning steps based on a novel first-token surprisal metric.
-Finally, ASAP teaches models to autonomously generate and leverage these concise CoTs at inference time, enabling efficient reasoning in coding tasks.
-Experiments show that ASAP achieves state-of-the-art accuracy across multiple code generation benchmarks while substantially reducing training and inference costs.
-On the challenging LiveCodeBench v4_v5 benchmark, our approach reduces token generation by 23.5% and inference latency by 43.5% compared to the strongest baseline, while achieving a competitive accuracy of 36.19% in Pass@1.
-Our results highlight a promising direction for building powerful and efficient LRMs.
-[^1] Code and model available at https://github.com/Zengwh02/ASAP.
+Leveraging the insight that logical branching choices are concentrated at the onset of reasoning steps, it then enables logic-aware pruning by selecting logically essential reasoning steps based on a novel first-token surprisal metric.
+Finally, ASAP distills the models to autonomously generate and leverage these concise CoTs at inference time, enabling efficient reasoning.
+Experiments show that ASAP achieves state-of-the-art accuracy across multiple benchmarks while substantially reducing training and inference costs.
+Pruning the Unsurprising: Efficient LLM Reasoning via First-Token Surprisal
+Wenhao Zeng1   Yaoning Wang2   Chao Hu1   Yuling Shi1 Chengcheng Wan3   Hongyu Zhang4   Xiaodong Gu1 1Shanghai Jiao Tong University  2Fudan University 3East China Normal University  4Chongqing University {zengwh_cs, xiaodong.gu}@sjtu.edu.cn
 Figure 1: Illustration of CoT pruning by ASAP.
-The original CoT emitted by LRMs contains redundant branches (highlighted in red dashed boxes).
-ASAP prunes them in the first stage.
-It then identifies the first tokens of reasoning steps (marked in blue), whose surprisal scores are computed to guide the second-stage pruning.
+The Original CoT generated by LRMs exhibits two types of redundancy: (1) Structural Redundancy, such as digressive branches (highlighted in red dashed boxes), which are removed by our Stage 1 Anchor-guided pruning; and (2) Logical Redundancy within valid paths.
+ASAP addresses the latter in Stage 2 by computing the surprisal of the first tokens of reasoning steps (marked in blue) to identify and retain only the critical cognitive pivots.
 1 Introduction
-Recently, the emergence of Large Reasoning Models (LRMs), including OpenAI’s o1 (Jaech et al.
-2024) and DeepSeek-R1 (Guo et al.
-2025), has demonstrated remarkable capabilities across a wide range of domains.
-LRMs empower LLMs with the Chain-of-Thought (CoT) prompting paradigm (Wei et al.
-2022), which prompts models to “think step by step”.
-By generating longer reasoning traces during inference, LRMs are capable of handling complex code reasoning tasks (Shi et al.
-2024).
-Despite achieving significant performance gains, the current LRMs come at a heavy computational cost when the CoTs scale up.
-Long CoTs often contain substantial redundancy, particularly pronounced in code generation tasks, where LRMs frequently explore diverse strategies or engage in self-debugging (Bi et al.
-2024; Xie et al.
-2025; Wu et al.
-2025; Qu et al.
-2025).
-This not only increases inference and training costs but also suggests a counterintuitive insight: removing redundant reasoning traces may improve both efficiency and performance.
-To improve the efficiency of CoT reasoning, a growing body of research has emerged on efficient reasoning (Qu et al.
-2025).
-A prominent line of work is CoT compression for efficient model fine-tuning, which aims to distill concise yet effective reasoning patterns into the model.
-For example, TokenSkip (Xia et al.
-2025a) adapts general-purpose context compressors such as Selective Context (Li et al.
-2023) and LLMLingua-2 (Pan et al.
-2024).
-However, these token-level compression methods are fundamentally ill-suited for code reasoning as they damage the logic of an entire code block within the trace, even if only a variable name is removed.
-The compressed context makes it hard for models to learn effective reasoning patterns.
-A more promising direction is step-level pruning.
-For example, SPIRIT (Cui et al.
-2025) prunes the reasoning steps while preserving structural and semantic integrity.
-However, these approaches face a deeper challenge: How to reliably estimate the logical importance of a reasoning step.
-SPIRIT relies on perplexity (PPL), but we argue that PPL is an inadequate metric as it primarily reflects linguistic fluency rather than logical necessity and has shown limitations in long-context scenarios (Hu et al.
-2024; Fang et al.
-2024).
-To overcome these challenges, we propose ASAP: A novel framework for Anchor-guided, SurprisAl-based Pruning of CoTs.
-Our core insight, illustrated in Figure 1, is identifying and pruning redundant steps in CoTs while preserving a coherent logical path.
-In the first stage, ASAP leverages the inherent structural and logical nature of code to generate a step-by-step reasoning path.
-The path serves as an anchor for coarse-grained pruning, efficiently identifying and preserving the core reasoning backbone.
-In the second stage, it enables a fine-grained logic-aware pruning by selecting the first tokens of reasoning steps based on a new First-Token Surprisal metric.
-The resulting CoT retains only the steps that introduce novel and meaningful information.
-By fine-tuning on the pruned CoTs, we teach a large reasoning model to generate and leverage these compact reasoning patterns at inference time, enabling efficient reasoning in coding tasks.
-We validate our approach through extensive experiments on the DeepSeek-R1-Distill-Qwen-7B and DeepSeek-R1-Distill-Llama-8B (Guo et al.
-2025) across a suite of challenging code generation benchmarks, including HumanEval (Chen et al.
-2021), HumanEval+ (Liu et al.
-2023), LiveCodeBench (Jain et al.
-2024), and LeetCodeDataset (Xia et al.
-2025b).
-Our resulting model establishes a new state-of-the-art in the performance-efficiency trade-off.
+The emergence of Large Reasoning Models, including OpenAI’s o1 Jaech et al.
+(2024) and DeepSeek-R1 Guo et al.
+(2025), marks a paradigm shift in artificial intelligence.
+By scaling up Chain-of-Thought (CoT) reasoning Wei et al.
+(2022), these models demonstrate emergent capabilities in complex domains such as mathematics Sun et al.
+(2025), programming Shi et al.
+(2024); Yang et al.
+(2025b); Hu et al.
+(2025), and logical reasoning Liu et al.
+(2025); Zhang et al.
+(2025c, b).
+However, this performance comes at a prohibitive cost: reasoning traces often span thousands of tokens, introducing substantial latency and memory overhead.
+Crucially, these lengthy traces often contain substantial redundancy, such as over-explaining simple problems or superficially exploring multiple paths for complex ones Bi et al.
+(2024); Xie et al.
+(2025); Wu et al.
+(2025); Qu et al.
+(2025).
+For instance, the Original CoT in Figure 1 contains tangential branches (highlighted in red dashed boxes), such as exploring an alternative manual implementation that is subsequently rejected (“But that’s more code…”).
+Furthermore, the reasoning is punctuated by syntactic fillers that contribute little to the core logic.
+This observation raises a fundamental question: Can we identify and retain only the “cognitive pivots” of reasoning while discarding the redundancy?
+A growing body of research has emerged on CoT compression for efficient reasoning Qu et al.
+(2025).
+Token-level methods like TokenSkip Xia et al.
+(2025a) adapt general-purpose context compressors such as LLMLingua-2 Pan et al.
+(2024) to prune non-informative tokens.
+However, indiscriminate token removal risks disrupting the syntactic integrity of the reasoning chain.
+To address this, step-level pruning methods like SPIRIT Cui et al.
+(2025) trim entire reasoning steps, thereby preserving structural coherence.
+Nevertheless, these approaches face a fundamental challenge: accurately estimating the logical importance of each step.
+They typically rely on fixed metrics like perplexity (PPL), which measures the overall predictability of a sentence.
+This holistic measure often dilutes the signal of critical logical leaps with the noise of syntactically predictable but logically trivial content.
+In this work, we ground CoT compression from an information-theoretic perspective.
+Through an empirical analysis of 10 million reasoning tokens (detailed in Section 2), we find that the logical progression within a CoT sequence is not uniformly distributed; instead, its information density is highly concentrated at the beginning of each reasoning step—specifically within the first few tokens (blue-highlighted in Figure 1).
+These tokens, such as “But” (self-correction) or “So” (deduction, not continuation), serve as high-entropy cognitive pivots.
+By leveraging the surprisal of these initial tokens, we can distinguish between critical logical transitions and predictable elaborations.
+Guided by this insight, we propose ASAP (Anchor-guided, SurprisAl-based Pruning), a coarse-to-fine framework designed to preserve these high-information steps.
+ASAP in a two-stage cascade that directly addresses the two types of redundancies identified in our study (illustrated in Figure 1): First, it employs Anchor-guided Pruning to remove structural redundancies.
+By generating a concise step-by-step reasoning trace as a logical backbone, it identifies and prunes the irrelevant branches (e.g., the red boxes in Figure 1).
+Second, it performs Surprisal-based Refining to eliminate logic-sparse steps.
+Leveraging our First-Token Surprisal metric, this stage iteratively filters out steps acting as mere fillers while retaining the high-surprisal cognitive pivots.
+Finally, we distill these compact, logic-dense CoTs into a target model, enabling it to generate efficient reasoning chains.
+We validate our approach through extensive experiments on the DeepSeek-R1-Distill-Qwen-7B and DeepSeek-R1-Distill-Llama-8B Guo et al.
+(2025) across diverse domains.
+The results demonstrate that ASAP establishes a superior Pareto frontier between performance and efficiency.
 Notably, on the challenging LiveCodeBench v4_v5 benchmark, ASAP achieves 36.19% Pass@1 while reducing token generation by 23.5% and inference latency by 43.5% compared to the strongest baseline.
 Our main contributions are summarized as follows:
-  • We propose ASAP, a novel two-stage pruning framework that combines coarse-grained anchor-guided pruning with fine-grained surprisal-based pruning to identify and retain logically critical steps.
-  • We are the first to propose and empirically validate First-Token Surprisal as a superior metric to perplexity for measuring the logical importance of CoT steps in the code generation task.
-  • Extensive experiments on multiple code generation benchmarks demonstrate that models fine-tuned on CoTs pruned by ASAP achieve state-of-the-art accuracy while substantially reducing both training and inference costs.
-2 Related Work
-2.1 Chain-of-Thought and Advanced Reasoning
-Chain-of-Thought (CoT) prompting has emerged as a foundational technique for eliciting reasoning in large language models (LLMs) (Wei et al.
-2022).
-This has sparked a surge of research exploring more sophisticated reasoning frameworks such as Tree of Thoughts (Yao et al.
-2023), Graph of Thought (Lei et al.
-2023), Natural Program (Ling et al.
-2023), etc.
-More recently, the emergence of powerful Large Reasoning Models (LRMs) such as OpenAI’s o1 (Jaech et al.
-2024) and DeepSeek-R1 (Guo et al.
-2025) has further advanced the field.
-These models exhibit strong long-chain reasoning capabilities, which are often enhanced through supervised fine-tuning (SFT) and reinforcement learning (RL) (Jaech et al.
-2024; Guo et al.
-2025; Kimi et al.
-2025; Yang et al.
-2025; Yu et al.
-2025; Wang et al.
-2025), especially in domains such as mathematical and programming reasoning.
-Unlike prior work that focuses on enhancing reasoning by generating longer or more complex CoTs, we aim to improve both reasoning performance and efficiency by pruning redundant steps from existing CoTs.
+  • We present an empirical analysis of the information concentration of CoTs, uncovering that the surprisal of the starting token for each CoT step is a more robust indicator of logical importance than perplexity.
+  • We propose ASAP, a novel CoT compression framework that combines structural alignment with information-guided refinement.
+  • Extensive experiments on multiple benchmarks demonstrate that models fine-tuned on CoTs pruned by ASAP achieve state-of-the-art accuracy while substantially reducing computational costs.
+(a) Entropy Distribution
+(b) High-Frequency First Tokens
+(c) High-Entropy Tokens
+Figure 2: Empirical analysis of 10M tokens from DeepSeek-R1-Distill-Qwen-32B.
+(a) The entropy distribution reveals a clear information concentration: first tokens (blue) exhibit significantly higher uncertainty (entropy) compared to body tokens (orange), which are highly deterministic.
+(b) The most frequent first tokens are a mixture of logical operators (e.g., Wait) and ubiquitous syntactic connectors (e.g., So).
+(c) High-entropy states filter out predictable fillers like So or Then, while exclusively highlighting cognitive pivots such as Perhaps, What, and Alternative.
+2 Empirical Analysis
+To investigate the intrinsic distribution of logical information in CoTs, we conducted a large-scale analysis on 10 million tokens generated by DeepSeek-R1-Distill-Qwen-32B Guo et al.
+(2025) across diverse reasoning benchmarks (AIME and LiveCodeBench Jain et al.
+(2024)).
+Information Concentration in CoTs.
+We analyze the entropy distribution of the first token of each reasoning step compared to all subsequent tokens.
+Entropy, in this context, quantifies the model’s uncertainty regarding the next state transition Shannon (1948); Malinin and Gales (2020); Kuhn et al.
+(2023); Wang et al.
+(2025); Cheng et al.
+(2025).
+As illustrated in Figure 2(a), a distinct concentration is observed.
+Starting tokens (blue) exhibit a dispersed distribution with a significantly higher 90th percentile.
+In contrast, other tokens (orange) are heavily concentrated near zero entropy, indicating that once a reasoning step is initiated, its subsequent elaboration is largely deterministic and syntactically driven.
+This empirical evidence confirms that the logical branching points, where the model actively deliberates on the reasoning path, are structurally concentrated at the beginning of each step.
+Identifying Cognitive Pivots with Entropy.
+Having identified the informative start tokens, we perform a more in-depth analysis of real logical pivots among the start tokens.
+We aim to distinguish between superficial syntactic connectors and real logical pivots.
+Figure 2(b) presents high-frequency start tokens that appear in CoTs, which mix connectors (“So”, “Let”) with reasoning markers (“Wait”, “But”).
+However, when focusing on tokens generated in high-entropy states (Figure 2(c)), a qualitative shift emerges.
+Predictable connectors like “So” and “Then” are effectively suppressed due to their low uncertainty.
+Instead, the distribution is dominated by terms representing cognitive pivots and state transitions, such as: 1) Exploration: “Alternative”, “Another” (proposing hypotheses or new paths).
+2) Causality: “Because”, “Since” (providing formal justification).
+3) Self-Correction: “Perhaps”, “What” (indicating error detection or logic reversal).
+This analysis demonstrates that high entropy is a robust indicator of logical salience.
+Since the actual next token is known in the given training sequence, we operationalize this insight by using First-Token Surprisal as a proxy to identify and preserve these critical reasoning hops in our proposed framework Fu et al.
+(2025).
 Figure 3: The overall framework of ASAP.
-The process begins with a training sample consisting of a question, an intermediate CoT, and the answer.
-In Stage 1, an LLM generates a “Direct CoT” ($C_{direct}$) from the question and answer.
-$C_{direct}$ acts as an anchor to prune the “Original CoT” ($C_{origin}$) into a “Coarse-grained Pruned CoT” ($C_{coarse}$).
-In Stage 2, we compute the first-token surprisal for each step in $C_{coarse}$ to remove less critical steps, yielding the final “Fine-grained Pruned CoT” ($C^{\prime}$) for fine-tuning.
-2.2 Context Compression for LLMs
-As tasks become more complex, the context in LLMs has expanded dramatically, resulting in significant computational overhead.
-This has motivated a line of work on context compression (Li et al.
-2025; Fang et al.
-2025), aiming to reduce input length while preserving key information.
-For instance, Selective Context (Li et al.
-2023) leverages self-information to delete redundant lexical units.
-LLMLingua (Jiang et al.
-2023) employs an LLM to select key demonstrations and then applies token-level filtering across the context.
-Its extension, LongLLMLingua (Jiang et al.
-2024), introduces document reordering and subsequence reconstruction to enable compression over longer contexts.
-LLMLingua-2 (Pan et al.
-2024) further advances this line by distilling training data and training a token classifier for task-agnostic compression.
-While these methods have proven effective for text compression, they are not tailored to the structured and logically coherent nature of CoT reasoning.
-In contrast, our method is explicitly designed for CoT compression, which identifies and preserves logically essential reasoning steps.
-2.3 Efficient Reasoning via Fine-Tuning
-The increasingly lengthy reasoning traces impose substantial costs, which have motivated a growing body of work focused on efficient reasoning (Qu et al.
-2025).
-Several recent approaches aim to compress CoTs while preserving their effectiveness.
-C3oT (Kang et al.
-2025) uses GPT-4 as a high-quality external compressor to generate shorter CoTs that retain essential reasoning content.
-TokenSkip (Xia et al.
-2025a) enables controllable compression by identifying and removing less important tokens.
-SPIRIT (Cui et al.
-2025) identifies important reasoning steps by detecting perplexity shifts when individual steps are removed.
-Another paradigm involves compressing CoTs into continuous latent representations, as explored by Coconut (Hao et al.
-2024), CCoT (Cheng and Van Durme 2024), CODI (Shen et al.
-2025), etc.
-Our method, ASAP, distinguishes itself from these methods through its explicit two-stage pruning framework, which first generates an anchor for coarse-grained pruning, followed by a fine-grained refinement based on first-token surprisal.
-By explicitly pruning at the step level, ASAP produces logically condensed human-readable CoTs.
+The pipeline consists of three phases: (1) In Stage 1, the LLM generates a “Direct Thought" ($\mathcal{P}$) from the (Question, Answer) pair.
+$\mathcal{P}$ acts as an anchor to prune the “Original CoT" ($C$) into a “Coarse-grained Pruned CoT" ($C_{coarse}$).
+(2) In Stage 2, we compute the First-Token Surprisal for each step in $C_{coarse}$.
+High-surprisal steps are retained, while low-surprisal fillers are pruned, yielding the final “Fine-grained Pruned CoT" ($C^{\prime}$).
+(3) In Training Stage, the data with ASAP pruned CoTs is used to fine-tune the LRM for efficient inference.
 3 Methodology
 3.1 Overall Framework
-Given a training corpus of triples $\{(Q_{i},\mathbf{C}_{i},A_{i})\}_{i=1}^{N}$—where each $Q_{i}$ denotes a question, $\mathbf{C}_{i}=\{s_{1},\dots,s_{N}\}$ is the intermediate reasoning steps, known as chain-of-thoughts (CoT), generated by a large reasoning model $\mathcal{L}$, and $A_{i}$ represents the predicted answer to the question.
-Our goal is to compress each $\mathbf{C}_{i}$ into a concise one $\mathbf{C}_{i}^{\prime}$ such that $|\mathbf{C}_{i}^{\prime}|\ll|\mathbf{C}_{i}|$ while the LRM maintains the quality of generated reasoning steps and answers when fine-tuned on the compressed triplets $\{Q_{i},\mathbf{C}_{i}^{\prime},A_{i}\}_{i=1}^{N}$.
-We propose ASAP, a novel method for CoT compression, as illustrated in Figure 2.
-ASAP adopts a coarse-to-fine pruning framework that firstly removes redundant steps from a given reasoning trace and then refines it by selecting the most critical logical steps.
-In the first stage called Anchor-guided Pruning, ASAP infers a concise, anchor reasoning path called “Direct Thoughts” ($C_{direct}$) based on the $(Q,A)$ pairs.
-Guided by $C_{direct}$ as an anchor, ASAP prunes the original CoT $C$ into a “coarse-grained pruned CoT” ($C_{coarse}$) using an LLM, which retains the main logical backbone while preliminarily removing digressions and redundancies.
-In the second stage called Surprisal-based Refining, ASAP performs a logic-aware refinement to $C_{coarse}$.
-A new metric called First-Token Surprisal is designed to measure the logical importance of a reasoning step and then iteratively filter steps with low surprisal, yielding the final CoT (denoted as $\mathbf{C}^{\prime}$ in our formulation).
-Finally, all $Q,C^{\prime},A$} tuplets are used to fine-tune a large reasoning model for downstream tasks, which internalizes concise reasoning patterns and substantially improves inference efficiency.
+Formally, we consider a supervised reasoning task defined by a dataset $\mathcal{D}=\{(Q_{i},C_{i},A_{i})\}_{i=1}^{N}$, where $Q_{i}$ is the query, $A_{i}$ is the predicted answer, and $C_{i}$ represents the original CoT generated by the LRMs.
+$C_{i}$ is a sequence of reasoning steps $C_{i}=\{s_{1},s_{2},\dots,s_{L}\}$.
+Our goal is to compress each $C_{i}$ into a concise pruned one $C_{i}^{\prime}$ such that $|C_{i}^{\prime}|\ll|C_{i}|$ while the LRMs maintains the quality of generated reasoning steps and answers when fine-tuned on the dataset $\mathcal{D^{\prime}}=\{(Q_{i},C_{i}^{\prime},A_{i})\}_{i=1}^{N}$.
+We propose ASAP, a coarse-to-fine framework tailored to the redundancy of “Original CoT" ($C$), as illustrated in Figure 3.
+Stage 1 (Anchor-guided Pruning) reduces structural redundancy (e.g., dead ends) by aligning the CoT with a generated logical backbone.
+The LLM generates a “Direct Thought" ($\mathcal{P}$) from the $(Q,A)$ pairs.
+$\mathcal{P}$ acts as an anchor to prune the $C$ into a “Coarse-grained Pruned CoT" ($C_{coarse}$).
+Stage 2 (Surprisal-based Refining) reduces logical redundancy (e.g., syntactic fillers) by filtering non-informative steps.
+We approximate the information of each step in $C_{coarse}$ with their surprisal of start tokens and prune low-surprisal steps, yielding the final “Fine-grained Pruned CoT" ($C^{\prime}$).
+Finally, all $\{(Q,C^{\prime},A)\}$ are utilized to fine-tune the target model.
 3.2 Anchor-guided Pruning
-Generating the Direct CoT Anchor
-We prompt an LLM to generate a concise and coherent anchor reasoning trace, $C_{direct}$, based on the question and answer.
-An anchor is defined as a structured, step-by-step explanation that outlines how to derive the answer from the question.
-This leverages the inherent logical and procedural nature of code reasoning.
-The resulting $C_{direct}$ serves as a logical backbone—an anchor—to guide the subsequent pruning of the more elaborate $C_{origin}$.
-The prompt used for the generation is shown in Appendix E.
-Pruning with Pattern Matching
-Using $C_{direct}$ as a reference, we prompt the LLM to prune the original CoT $C$.
-Specifically, the LLM is instructed to: 1) remove unnecessary reasoning steps from $C$; 2) retain all key supporting content that aligns with the logic of $C_{direct}$; and 3) crucially, preserve the original wording without introducing new information.
-The prompt used for pruning is shown in Appendix E.
-To mitigate LLM hallucination, we incorporate a validation step based on Gestalt Pattern Matching (Black 2004), which validates structural and semantic alignment with $C$.
+Directly pruning raw CoTs is challenging due to the noise and unstructured digressions inherent in LLM reasoning Zhou et al.
+(2024).
+To address this, we first construct a high-level logical skeleton to narrow the pruning space.
+Generate Direct Thoughts.
+We prompt the LLM to infer a concise reasoning path called “Direct Thought” ($\mathcal{P}$) based on the $(Q,A)$ pair (see Appendix B for prompts).
+Unlike exploratory CoTs, $\mathcal{P}$ is generated as a structured, step-by-step explanation that outlines how to derive the answer from the question, exemplified in Appendix B.
+This $\mathcal{P}$ acts as a reference anchor, outlining the least reasoning trajectory required to solve the problem.
+Pruning with Pattern Matching.
+Guided by the anchor $\mathcal{P}$, we prompt the LLM to prune the original CoT $C$.
+Specifically, the LLM is instructed to: 1) remove unnecessary reasoning steps from $C$; 2) retain all key supporting content that aligns with the logic of $\mathcal{P}$; and 3) crucially, preserve the original wording without introducing new information.
+The prompt used for pruning is shown in Appendix B.
+The goal is to extract the subsequence of $C$ that semantically aligns with $\mathcal{P}$ while discarding irrelevant branches (as shown in the “Original CoT” block of Figure 3).
+Crucially, to mitigate LLM hallucination during compression, we enforce an extractive constraint, which validates structural and semantic alignment with $C$.
 Specifically, we design a pattern-matching algorithm that verifies whether each step in $C_{coarse}$ corresponds to a matching step in $C$ while preserving their original order.
-The matching is performed using Gestalt Pattern Matching (Black 2004) as a text similarity metric.
+The matching is performed using Gestalt Pattern Matching Black (2004) as a text similarity metric.
 A pruning is considered valid only if all steps in $C_{coarse}$ achieve a similarity score above a predefined threshold $\tau$ when matched against sequential steps in $C$.
-This guarantees that $C_{coarse}$ forms a consistent substructure of the original reasoning trace.
-The full pattern-matching algorithm is detailed in Algorithm 1.
+The full pattern-matching algorithm is detailed in Algorithm 1 (see Appendix A).
 We leverage high-temperature sampling, which provides the necessary diversity to efficiently re-prompt failed cases, ensuring that a valid $C_{\text{coarse}}$ can be eventually generated.
+3.3 Surprisal-based Refining
+Following the coarse-grained pruning, the resulting $C_{coarse}$ may still contain verbose steps that contribute little to the logic.
+Grounded in our empirical finding that logical information is concentrated (Section 2), we perform a meticulous, logic-aware refinement in $C_{coarse}$ to identify more subtle redundancies within the core reasoning path.
+First-Token Surprisal as Logical Importance.
+We introduce First-Token Surprisal as a novel metric to precisely quantify the logical importance of each step, enabling us to filter out the least informative ones and produce the final highly condensed CoT.
+Let a reasoning step $s$ be a sequence of tokens $s=(x_{1},x_{2},\dots,x_{T})$.
+The informational value of $s$ within the context of previous steps $\mathcal{C}_{pre}$ is typically estimated by its joint probability.
+However, our analysis reveals that the first token $x_{1}$ serves as the “cognitive pivot” carrying the majority of the uncertainty.
+Therefore, we define the First-Token Surprisal $\mathcal{S}(s)$ as:
+$\mathcal{S}(s\mid\mathcal{C}_{pre})=-\log P_{\theta}(x_{1}\mid\mathcal{C}_{pre})$ (1)
+where $P_{\theta}$ denotes the probability distribution of the LRM.
+A high $\mathcal{S}(s)$ indicates a high-information transition (e.g., initiating a new deduction or self-correction), whereas a low score suggests a deterministic continuation or syntactic filler.
+Pruning using First-Token Surprisal.
+We formulate the fine-grained pruning as a constrained maximization problem.
+Our goal is to select a subset of steps $S^{\prime}\subset C_{coarse}$ that maximizes the total logical information subject to a length budget $L_{max}$:
+$\begin{split}S^{*}=\operatorname*{arg\,max}_{S^{\prime}\subseteq C_{coarse}}\sum_{s\in S^{\prime}}\mathcal{S}(s)\\ \text{s.t.}\quad\sum_{s\in S^{\prime}}\text{len}(s)\leq L_{max}\end{split}$ (2)
+This formulation explicitly prioritizes steps with high information density.
+To solve this efficiently, we employ a greedy iterative strategy.
+We calculate the surprisal score for all steps in $C_{coarse}$ and iteratively remove the step with the lowest $\mathcal{S}(s)$, while the relative order of steps in $S^{\prime}$ is preserved.
+The detailed procedure is provided in Algorithm 2 (see Appendix A).
+This process yields the final fine-grained CoT $C^{\prime}$, which retains the critical “aha moments” Guo et al.
+(2025) while meeting efficiency constraints.
+3.4 Supervised Fine-tuning
+Following the pruning, we construct the final training dataset $\mathcal{D}^{\prime}=\{(Q_{i},\mathbf{C}_{i}^{\prime},A_{i})\}_{i=1}^{N}$.
+For each instance, we concatenate the pruned CoT ($\mathbf{C}_{i}^{\prime}$) and the final answer ($A_{i}$) to form the complete target response $R_{i}$.
+We then fine-tune the LRM to minimize the standard negative log-likelihood of the target response tokens, conditioned on the input question.
+Formally, the loss is defined as:
+$\mathcal{L}=-\sum_{i=1}^{N}\sum_{j=1}^{|R_{i}|}\log P_{\theta}(r_{i,j}|Q_{i},r_{i,<j})$ (3)
+where $r_{i,j}$ is the $j$-th token of the target response $R_{i}$, and $\theta$ represents the parameters of the model being fine-tuned.
+This supervised fine-tuning process effectively distills the knowledge from our pruning framework into the model.
+By training on these compact, logically salient examples, the model learns to internalize efficient reasoning patterns.
+4 Experiments
+Methods HE+ LCBv1_v3 LCBv4_v5 LCD
+Acc $\uparrow$ Tok $\downarrow$ Lat $\downarrow$ Acc $\uparrow$ Tok $\downarrow$ Lat $\downarrow$ Acc $\uparrow$ Tok $\downarrow$ Lat $\downarrow$ Acc $\uparrow$ Tok $\downarrow$ Lat $\downarrow$
+Zero-shot 68.29 3051 1.16 42.16 7088 3.59 25.37 8336 5.15 19.74 8680 4.95
+Original 75.61 2973 1.12 52.12 6611 3.15 30.97 8289 4.83 25.00 8485 4.72
+Selective Context 54.88 2979 1.13 30.23 7025 3.75 16.79 8558 5.35 15.79 8461 4.90
+LLMLingua-2 68.29 3075 1.19 38.89 6953 3.60 22.76 8474 5.31 17.54 8513 4.81
+TokenSkip 73.78 2823 1.07 32.35 7095 3.85 20.15 8400 5.37 18.42 8503 4.87
+SPIRIT 75.61 2764 1.07 50.82 6524 3.09 33.58 7892 4.62 25.00 8186 4.45
+ASAP 78.66 2464 0.98 54.74 5177 2.09 36.19 6035 2.61 27.63 7541 3.48
+Table 1: Experimental results of different methods on code generation benchmarks with DeepSeek-R1-Distill-Qwen-7B.
+We report accuracy (Acc), average number of generated tokens (Tok), and average generation latency (Lat) measured in seconds.
+The best results are highlighted in bold, and the second-best are underlined.
+Methods GSM8K MATH500 AIME24 AIME25
+Acc $\uparrow$ Tok $\downarrow$ Lat $\downarrow$ Acc $\uparrow$ Tok $\downarrow$ Lat $\downarrow$ Acc $\uparrow$ Tok $\downarrow$ Lat $\downarrow$ Acc $\uparrow$ Tok $\downarrow$ Lat $\downarrow$
+Zero-shot 83.55 1301 0.27 60.40 2629 0.70 36.67 8352 6.76 40.00 8145 6.67
+Original 86.35 1250 0.26 63.80 2511 0.66 46.67 8034 6.43 43.33 8026 6.75
+Selective Context 75.44 1108 0.24 52.20 2507 0.66 16.67 9610 7.40 10.00 9329 7.40
+LLMLingua-2 79.98 1128 0.24 54.60 2802 0.76 36.67 8369 6.60 23.33 8919 7.42
+TokenSkip 85.37 1303 0.27 65.60 2483 0.65 36.67 8073 6.61 33.33 8465 7.44
+SPIRIT 88.55 1118 0.23 64.20 2144 0.57 46.67 7198 5.78 43.33 7817 6.57
+ASAP 90.75 753 0.16 70.80 1649 0.43 46.67 5552 5.04 36.67 5434 5.10
+Table 2: Experimental results of different methods on mathematical reasoning benchmarks with DeepSeek-R1-Distill-Qwen-7B.
+We report accuracy (Acc), average number of generated tokens (Tok), and average generation latency (Lat) measured in seconds.
+The best results are highlighted in bold, and the second-best are underlined.
+4.1 Experimental Setup
+Models and Datasets.
+All experiments are conducted on the DeepSeek-R1-Distill-Qwen-7B and DeepSeek-R1-Distill-Llama-8B Guo et al.
+(2025), with DeepSeek-R1-Distill-Qwen-7B as the default backbone across all settings.
+For the code reasoning domains, we use the Python subset of the CodeForces-CoTs Hugging Face (2025) dataset.
+For the math reasoning domain, we adopt the OpenR1-Math Hugging Face (2025) dataset and randomly sample 10K instances to match the size of the code subset, ensuring a balanced comparison across domains.
+The datasets consist of high-quality Chain-of-Thought (CoT) samples generated by DeepSeek-R1, making it particularly suitable for training competitive reasoning tasks.
+Detailed implementation settings (hyperparameters, hardware, etc.) are provided in Appendix C.
+Benchmarks.
+We evaluate our method on a suite of widely used benchmarks that cover both code generation and mathematical reasoning tasks.
+For code generation, we adopt HumanEval+ Chen et al.
+(2021); Liu et al.
+(2023), LiveCodeBench v1_v3, LiveCodeBench v4_v5 Jain et al.
+(2024), and LeetCodeDataset Xia et al.
+(2025b).
+For mathematical reasoning, we evaluate on GSM8K Cobbe et al.
+(2021), MATH500 Hendrycks et al.
+(2021), AIME24, and AIME25.
+Baselines.
+We compare our method against a comprehensive set of baselines.
+Zero-shot refers to the original model without any task-specific fine-tuning.
+Original denotes the model fine-tuned on the uncompressed CoTs from the training data.
+Among compression approaches, Selective Context Li et al.
+(2023) prunes redundant lexical units based on self-information; LLMLingua-2 Pan et al.
+(2024) distills GPT-4’s token importance signals into a lightweight Transformer encoder trained as a token classifier; TokenSkip Xia et al.
+(2025a) learns to skip less informative tokens to achieve controllable compression; and SPIRIT Cui et al.
+(2025) identifies critical reasoning steps by measuring perplexity shifts.
+Except for the zero-shot setting, all methods involve fine-tuning on CoTs processed according to their respective compression strategies.
+Metrics.
+We evaluate both accuracy and inference efficiency of each approach across three metrics: Pass@1 (Acc), which measures the percentage of problems correctly solved on the first attempt; Tokens (Tok), which denotes the average number of tokens generated by the LRMs; and Latency (Lat), which measures the average time (in seconds) required for the model generation.
+4.2 Main Results
+Tables 1 and 2 present the results of various methods on all benchmarks.
+The results show that the model fine-tuned on CoTs pruned by ASAP consistently achieves the best trade-off between accuracy and efficiency.
+It achieves the best accuracy while generating the fewest tokens, leading to the lowest generation latency.
+We notice a clear distinction between token-level and step-level pruning strategies.
+Token-level baselines such as Selective Context, LLMLingua-2, and TokenSkip exhibit a significant performance degradation compared to the original CoTs.
+This is because the token removal disrupts the syntactic structure and semantic coherence of the original reasoning steps.
+Consequently, the fine-tuning data becomes fragmented and grammatically unnatural, making it difficult for the model to learn the intended logical flow of the CoT.
+Step-level methods, such as SPIRIT, perform significantly better than token-level pruning methods, due to the preservation of sentence-level integrity.
+While SPIRIT improves efficiency over the Original with comparable accuracy, our method achieves higher efficiency and accuracy at the same time.
+This improvement is particularly pronounced on the challenging LiveCodeBench v4_v5 benchmark: ASAP reduces the average number of generated tokens by 23.5% (from 7892 to 6035) and lowers generation latency by 43.5% (from 4.62s to 2.61s), while also achieving a 7.8% improvement in accuracy (Pass@1 increases from 33.58% to 36.19%).
+4.3 Ablation and Analysis
+Effect of Different Components.
+To validate the contribution of each component, we conduct an ablation study on three model variants.
+1) w/o Anchor-guided Pruning: which skips Stage 1 and applies only surprisal-based pruning to the original CoT.
+2) w/o Surprisal-based Refining: which omits the surprisal-based refinement stage; and 3) w/o Both Pruning: equivalent to the original baseline, where the model is fine-tuned on the full, uncompressed CoT.
+Table 3 presents the results on LiveCodeBench v4_v5, which is representative of the consistent trends observed across benchmarks.
+Additional results are included in Appendix D.
+The results show that both pruning stages are essential and mutually complementary for optimal accuracy and efficiency.
+First, removing the anchor-guided pruning leads to a drop in both accuracy and efficiency.
+While the accuracy decrease is modest, the generation latency increases by a substantial 76.2% (from 2.61s to 4.60s), underscoring the importance of stage 1.
+Second, removing the surprisal-based refining results in a significant degradation across all metrics.
+The accuracy drops by 12.4% (Pass@1 decreases from 36.19% to 31.72%) relative to the ASAP, and efficiency improvements are largely lost.
+This highlights that our surprisal-based pruning mechanism is essential to select the most critical steps.
+Variants Acc $\uparrow$ Tok $\downarrow$ Lat $\downarrow$
+ASAP 36.19 6035 2.61
+w/o Anchor-guided Pruning 35.07 7735 4.60
+w/o Surprisal-based Refining 31.72 8061 4.83
+w/o Both Pruning 30.97 8289 4.83
+Table 3: Ablation study of different pruning strategies on LiveCodeBench v4_v5.
+We report accuracy (Acc), average number of generated tokens (Tok), and average generation latency (Lat) measured in seconds.
+Generalization to Different Architectures.
+To validate the generalizability of ASAP, we replicate our main experiments on the DeepSeek-R1-Distill-Llama-8B.
+Following the same experimental protocol, we compare ASAP against three strong baselines: Zero-shot, Original, and SPIRIT.
+We observe consistent trends across all benchmarks.
+For brevity, we present representative results on two key benchmarks: LiveCodeBench v4_v5 and AIME24 in Table 4, while reporting the full results in the Appendix E.
+The results in the Llama3.1 series are highly consistent with our findings in the Qwen2.5 series, confirming the generalization of the ASAP.
+As shown in Table 4, ASAP achieves the highest accuracy on both benchmarks, and the efficiency improvements are even more pronounced.
+On LiveCodeBench, for instance, ASAP not only surpasses the accuracy of the Original baseline (32.84% vs.
+31.34%) but also generates 49.1% fewer tokens and reduces latency by over 3x (from 8.60s to 2.69s).
+This suggests that the ASAP is particularly effective in identifying and distilling the core reasoning patterns, validating its robustness and broad applicability for improving reasoning efficiency across different model families.
+Methods LCB AIME
+Acc $\uparrow$ Tok $\downarrow$ Lat $\downarrow$ Acc $\uparrow$ Tok $\downarrow$ Lat $\downarrow$
+Zero-shot 25.00 8508 8.90 33.33 8445 10.42
+Original 31.34 8202 8.60 36.67 8550 10.04
+SPIRIT 30.22 7913 8.45 36.67 8788 10.04
+ASAP 32.84 4175 2.69 36.67 5314 6.97
+Table 4: Experimental results of different methods with DeepSeek-R1-Distill-Llama-8B.
+We report accuracy (Acc), average number of generated tokens (Tok), and average generation latency (Lat) measured in seconds.
+The best results are highlighted in bold.
+Figure 4: Performance of ASAP on LiveCodeBench v4_v5 under different token budgets.
+Impact of Token Budget.
+To evaluate the scalability and resource sensitivity of our method, we analyze its behavior under varying inference-time token budgets (i.e., the maximum number of tokens to the model).
+We compare ASAP against the three strong baselines-SPIRIT, Original, and Zero-shot—across all benchmarks, and observe consistent trends.
+For clarity, we present results on LiveCodeBench v4_v5 under six budget settings ranging from 2K to 12K tokens.
+Results for other benchmarks and additional statistics are provided in Appendix F.
+As shown in Figure 4, ASAP consistently outperforms all baselines across all budget settings.
+In particular, ASAP exhibits smooth performance scaling with respect to the token budget.
+We note that ASAP achieves superior performance-efficiency trade-offs.
+For example, ASAP with just an 8K token budget achieves higher accuracy than SPIRIT and Original at a much larger 12K budget.
+These results further validate the practical utility of ASAP in real-world scenarios.
+Training Efficiency.
+To quantify the training efficiency gains, we present results of the CodeForces-CoTs dataset in Table 5 and results on other datasets are provided in Appendix G.
+The results highlight the training efficiency advantage of the ASAP.
+By generating the most compact yet logically rich CoTs, our approach significantly reduces training overhead.
+Compared to the uncompressed baseline (Original), our method reduces the number of training tokens by 75.6% and shortens training time by 60.7%.
+These savings substantially exceed those achieved by all other baselines.
+ASAP enables a more resource-efficient training process, making it a practical and cost-effective solution for real-world deployment.
+Methods Tokens Time
+Original 13023 80.11
+Selective Context 6722 (-48.4%) 63.41 (-20.9%)
+LLMLingua-2 6919 (-46.9%) 65.25 (-18.6%)
+TokenSkip 9813 (-24.6%) 77.27 (-3.6%)
+SPIRIT 6082 (-53.3%) 57.45 (-28.3%)
+ASAP 3178 (-75.6%) 31.48 (-60.7%)
+Table 5: Training efficiency comparison on CodeForces-CoTs dataset.
+We report the average number of tokens per sample and training time measured in seconds per step.
+Percentages indicate the reduction relative to the Original baseline.
+5 Related Work
+Chain-of-Thought and Advanced Reasoning.
+Chain-of-Thought (CoT) prompting Wei et al.
+(2022) has evolved from heuristic prompting strategies Yao et al.
+(2023); Lei et al.
+(2023); Ling et al.
+(2023) to the training of specialized Large Reasoning Models (LRMs) like OpenAI’s o1 Jaech et al.
+(2024) and DeepSeek-R1 Guo et al.
+(2025).
+These models leverage reinforcement learning to scale test-time compute, generating lengthy reasoning traces to solve complex tasks Kimi et al.
+(2025); Yang et al.
+(2025a); Yu et al.
+(2025); Wang et al.
+(2025); Zhang et al.
+(2025a, d).
+Unlike prior works that enhance performance by scaling up CoT length, we focus on pruning redundancy to improve efficiency without compromising reasoning performance.
+Context Compression for LLMs.
+To mitigate the computational cost of long contexts, various compression techniques have been proposed Zhang et al.
+(2025e).
+Approaches like Selective Context Li et al.
+(2023), LLMLingua series Jiang et al.
+(2023); Pan et al.
+(2024), and LongCodeZip Shi et al.
+(2025) employ information-theoretic metrics or small external models to filter redundant tokens.
+However, these methods typically treat input as unstructured text.
+Applying them directly to CoT often disrupts the syntactic and logical coherence required for valid reasoning, a limitation that our pruning aims to overcome.
+Efficient Reasoning via Fine-Tuning.
+Recent research has explored various efficiency mechanisms Qu et al.
+(2025), ranging from compressing thoughts into continuous latent representations Hao et al.
+(2024); Cheng and Van Durme (2024); Shen et al.
+(2025) to compressing CoTs Kang et al.
+(2025); Xia et al.
+(2025a); Cui et al.
+(2025).
+Approaches like TokenSkip Xia et al.
+(2025a) and SPIRIT Cui et al.
+(2025) reduce length by filtering tokens or steps based on heuristics or perplexity shifts.
+However, these metrics often struggle to differentiate between syntactic fluency and logical necessity.
+ASAP differs by combining anchor-guided structural pruning with first-token surprisal, offering a more robust proxy for cognitive pivots.
+6 Conclusion
+In this paper, we address the inefficiency of Large Reasoning Models stemming from the structural and logical redundancies in Chain-of-Thought reasoning.
+Grounded in an information-theoretic perspective, our large-scale empirical analysis reveals a fundamental property of reasoning traces: Information Concentration, where the logical uncertainty is highly concentrated at the onset of reasoning steps.
+Guided by this insight, we propose ASAP.
+This coarse-to-fine framework first aligns the reasoning structure with a logical anchor and then refines it using a novel First-Token Surprisal metric.
+Extensive experiments across multiple benchmarks demonstrate that ASAP outperforms existing baselines, establishing a new state-of-the-art Pareto frontier between accuracy and efficiency.
+Our work highlights the potential of using information-theoretic signals for efficient reasoning.
+Future work will explore applying ASAP to online inference acceleration.
+Limitations
+While ASAP demonstrates significant improvements in reasoning efficiency, we acknowledge several limitations.
+First, our method relies on the availability of a capable LLM to generate high-quality “Direct Thoughts” in Stage 1.
+If the anchor contains logical errors or hallucinations, it may misguide the subsequent pruning, although our pattern-matching constraint mitigates this risk.
+Second, our experiments primarily focus on code generation and mathematical reasoning.
+While we believe the principle of information concentration applies broadly, the effectiveness of ASAP on creative writing or commonsense reasoning tasks remains to be verified.
+Appendix A Algorithms
 Algorithm 1 Pattern Matching
@@
 24:end function
-3.3 Surprisal-based Refining
-Following the coarse-grained pruning, we perform a meticulous, logic-aware refinement in $C_{coarse}$ to identify more subtle redundancies within the core reasoning path.
-First-Token Surprisal as Logical Importance
-We introduce First-Token Surprisal as a novel metric to precisely quantify the logical importance of each step, enabling us to iteratively filter out the least informative ones and produce the final highly condensed CoT.
-In information theory, entropy is used to quantify the average level of uncertainty or information associated with a variable’s potential states or possible outcomes (Shannon 1948).
-In the era of LLMs, this concept can be applied to each token generated by the model.
-Given an LLM generates a sequence of tokens $x_{1},x_{2},\dots,x_{N}$, the conditional probability of the $n$-th token is given by $p(x_{n}|x_{<n})$, derived from the model’s logits.
-Token-level entropy measures the model’s uncertainty over the entire vocabulary $V$ for the next token (Malinin and Gales 2020; Kuhn, Gal, and Farquhar 2023).
-It is the surprisal over all possible outcomes:
-$H(x_{n}|x_{<n})=-\sum_{v\in V}p(v|x_{<n})\log p(v|x_{<n})$ (1)
-In large reasoning models, the first token of each step often signals a logical transition or continuation.
-High entropy indicates the model’s uncertainty, suggesting that the step may introduce semantically or logically significant information.
-Conversely, low entropy typically reflects predictable continuations or syntactic fillers, implying limited contribution to the overall reasoning process (Wang et al.
-2025; Cheng et al.
-2025).
-However, since the actual next token is known in the sequence, we directly compute the surprisal of the target token, a more efficient and targeted metric that quantifies the unexpectedness of the target token $x_{t}$:
-$S(x_{t}|x_{<n})=-\log p(x_{t}|x_{<n})$ (2)
-Surprisal quantifies how unexpected a given token is to the model.
-We focus specifically on the first-token surprisal of each reasoning step.
-A high surprisal value at the beginning of a step indicates logical salience—that is, the step either initiates a new line of reasoning or introduces novel information.
-In contrast, a low surprisal value suggests a continuation of prior context or redundant elaboration.
-Accordingly, we prune low-surprisal steps to obtain more concise and informative reasoning traces.
-Iterative Pruning Algorithm
-Based on the first-token surprisal metric, we propose a length-controllable iterative pruning algorithm.
-As shown in Algorithm 2, it calculates the surprisal score for the first token of each reasoning step within the $C_{coarse}$.
-It then iteratively removes the step with the lowest surprisal score until the total length of the CoT falls below a predefined token budget $L_{max}$.
-This process yields the final $C^{\prime}$, which is highly compressed while retaining the most logically significant steps.
 Algorithm 2 Iterative Pruning via First-Token Surprisal
@@
 22:end function
-3.4 Supervised Fine-tuning
-Following the two-stage pruning by our ASAP framework, we construct the final training dataset using the compressed triplets $\{(Q_{i},\mathbf{C}_{i}^{\prime},A_{i})\}_{i=1}^{N}$.
-For each instance, we concatenate the pruned CoT ($\mathbf{C}_{i}^{\prime}$) and the final answer ($A_{i}$) to form the complete target response, denoted as $R_{i}$.
-We then fine-tune a large reasoning model on this dataset.
-The training goal is to minimize the negative log-likelihood of the target response tokens, conditioned on the input question.
-Formally, the loss is defined as:
-$\mathcal{L}=-\sum_{i=1}^{N}\sum_{j=1}^{|R_{i}|}\log P_{\theta}(r_{i,j}|Q_{i},r_{i,<j})$ (3)
-where $r_{i,j}$ is the $j$-th token of the target response $R_{i}$, and $\theta$ represents the parameters of the model being fine-tuned.
-This supervised fine-tuning process effectively distills the knowledge from our pruning framework into the model.
-By training on these compact and logically salient examples, the model learns to internalize the efficient reasoning patterns.
-4 Experiments
-Methods HumanEval LiveCodeBench LeetCodeDataset
-HE HE+ v1_v3 v4_v5 All
-Acc $\uparrow$ Tok $\downarrow$ Lat $\downarrow$ Acc $\uparrow$ Tok $\downarrow$ Lat $\downarrow$ Acc $\uparrow$ Tok $\downarrow$ Lat $\downarrow$ Acc $\uparrow$ Tok $\downarrow$ Lat $\downarrow$ Acc $\uparrow$ Tok $\downarrow$ Lat $\downarrow$
-Zero-shot 73.78 3051 1.16 68.29 3051 1.16 42.16 7088 3.59 25.37 8336 5.15 19.74 8680 4.95
-Original 78.05 2973 1.12 75.61 2973 1.12 52.12 6611 3.15 30.97 8289 4.83 25.00 8485 4.72
-Selective Context 59.76 2979 1.13 54.88 2979 1.13 30.23 7025 3.75 16.79 8558 5.35 15.79 8461 4.90
-LLMLingua-2 71.34 3075 1.19 68.29 3075 1.19 38.89 6953 3.60 22.76 8474 5.31 17.54 8513 4.81
-TokenSkip 76.83 2823 1.07 73.78 2823 1.07 32.35 7095 3.85 20.15 8400 5.37 18.42 8503 4.87
-SPIRIT 83.54 2764 1.07 75.61 2764 1.07 50.82 6524 3.09 33.58 7892 4.62 25.00 8186 4.45
-ASAP 84.15 2464 0.98 78.66 2464 0.98 54.74 5177 2.09 36.19 6035 2.61 27.63 7541 3.48
-Table 1: Experimental results of different methods with DeepSeek-R1-Distill-Qwen-7B.
-We report accuracy (Acc), average number of generated tokens (Tok), and average generation latency (Lat) measured in seconds.
-The best results are highlighted in bold, and the second-best are underlined.
-4.1 Experimental Setup
-Models and Datasets
-All experiments are conducted on the DeepSeek-R1-Distill-Qwen-7B and DeepSeek-R1-Distill-Llama-8B (Guo et al.
-2025), with DeepSeek-R1-Distill-Qwen-7B as the default backbone across all settings.
-We use the Python subset of the CodeForces-CoTs (Hugging Face 2025) dataset for training.
-The dataset consists of high-quality Chain-of-Thought (CoT) samples generated by DeepSeek-R1, making it particularly suitable as training data for competitive programming reasoning tasks.
-Benchmarks
-We evaluate our method on a suite of widely used code generation benchmarks that span a range of difficulty levels:
-  • HumanEval (Chen et al. 2021) consists of 164 hand-crafted programming problems. HumanEval+ (Liu et al. 2023) extends HumanEval with approximately 80 times more unique test cases.
-  • LiveCodeBench (Jain et al.
-2024) is a comprehensive, contamination-free benchmark that continuously aggregates new problems from competitive programming platforms.
-We use two subsets: v1_v3 (612 problems) and v4_v5 (268 problems).
-  • LeetCodeDataset (Xia et al.
-2025b) provides a high-quality evaluation set of 228 problems collected from LeetCode.
-It covers over 90% of Python problems on the platform, with each problem featuring over 100 test cases.
-Baselines
-We compare our method against a comprehensive set of baselines:
-  • Zero-shot: The original model without any task-specific fine-tuning.
-  • Original: The model is fine-tuned on the original, uncompressed CoTs from the training data.
-  • Selective Context (Li et al. 2023): Identifies and prunes redundant lexical units in the input using self-information, aiming to retain only the most informative tokens.
-  • LLMLingua-2 (Pan et al.
-2024): Distills GPT-4’s token importance signals into a lightweight Transformer encoder, which is trained as a token classifier.
-This approach compresses content using the token classifier.
-  • TokenSkip (Xia et al. 2025a): Trains a model to selectively skip less important tokens in a CoT, allowing for controllable compression based on token importance scores.
-  • SPIRIT (Cui et al. 2025): Identifies critical reasoning steps by measuring perplexity shifts—steps are retained if their removal significantly increases the perplexity of the remaining text.
-Except for the zero-shot baseline, all other methods involve fine-tuning the model using CoTs that have been processed according to their strategies.
-Metrics
-To comprehensively evaluate each method, we report both accuracy and inference efficiency:
-  • Accuracy is measured by Pass@1, which denotes the percentage of problems correctly solved on the first attempt.
-  • Efficiency is evaluated using two metrics: the average number of generated tokens and the average generation latency measured in seconds per sample.
-Implementation Details
-All models are fully fine-tuned (i.e., full-parameter fine-tuning) using Unsloth (Daniel Han and team 2023) for efficiency.
-For inference, we utilize the vLLM (Kwon et al.
-2023) engine to enable fast and efficient decoding.
-To ensure reproducibility, we use greedy decoding and disable prefix caching.
-The default generation token budget is adjusted based on task difficulty: 6K for HumanEval and HumanEval+, and 10K for LiveCodeBench and LeetCodeDataset.
-All experiments are conducted on NVIDIA H20 GPUs.
-Additional implementation details are provided in Appendix A.
-4.2 Main Results
-The comprehensive results of our method compared to all baselines are presented in Table 1.
-The results show that the model fine-tuned on CoTs pruned by ASAP consistently achieves the best trade-off between accuracy and efficiency.
-It outperforms all baselines in terms of accuracy, while generating the fewest tokens and achieving the lowest generation latency across all benchmarks.
-Statistical analysis using the Wilcoxon signed-rank test confirms that the improvements of our method over all baselines are statistically significant ($p<0.05$).
-A deeper analysis of the results reveals a clear distinction between token-level and step-level pruning strategies.
-Token-level baselines like Selective Context, LLMLingua-2, and TokenSkip exhibit a significant performance degradation compared to the original CoTs.
-This is because the token removal disrupts the syntactic structure and semantic coherence of the original reasoning steps.
-Consequently, the fine-tuning data becomes fragmented and grammatically unnatural, making it difficult for the model to learn the intended logical flow of the CoT.
-Step-level methods, such as SPIRIT, perform significantly better than token-level pruning methods, due to the preservation of sentence-level integrity.
-While SPIRIT improves efficiency over the Original with comparable accuracy, our method achieves higher efficiency and accuracy at the same time.
-This improvement is particularly pronounced on the challenging LiveCodeBench v4_v5 benchmark: ASAP reduces the average number of generated tokens by 23.5% (from 7892 to 6035) and lowers generation latency by 43.5% (from 4.62s to 2.61s), while also achieving a 7.8% improvement in accuracy (Pass@1 increases from 33.58% to 36.19%).
-These results verify our core hypothesis that pruning reasoning chains based on first-token surprisal is more effective than perplexity or token-level pruning.
-4.3 Ablation Study
-To validate the contribution and necessity of each component in our two-stage pruning framework, we conduct a detailed ablation study. Specifically, we evaluate the following three variants:
-  • ASAP w/o Coarse-grained Pruning: This variant skips Stage 1 and applies only the fine-grained surprisal-based pruning directly to the original CoT.
-  • ASAP w/o Fine-grained Pruning: This variant applies only the coarse-grained anchor-guided pruning from Stage 1, omitting the surprisal-based refinement in Stage 2.
-  • ASAP w/o Any Pruning: Equivalent to the Original baseline, where the model is fine-tuned on the full, uncompressed CoT without any pruning.
-Table 2 reports results on LiveCodeBench v4_v5, which is representative of the consistent trends observed across benchmarks. Additional results are included in Appendix B.
-The ablation study demonstrates that both pruning stages are essential and mutually complementary for achieving optimal accuracy and efficiency.
-First, removing the coarse-grained pruning stage (w/o Coarse-grained Pruning) leads to a drop in both accuracy and efficiency.
-While the accuracy decrease is modest, the generation latency increases by a substantial 76.2% (from 2.61s to 4.60s), underscoring the importance of stage 1 in eliminating redundancy and narrowing the search space for subsequent processing.
-Second, removing the fine-grained pruning stage (w/o Fine-grained Pruning) results in a significant degradation across all metrics.
-The accuracy drops by 12.4% (Pass@1 decreases from 36.19% to 31.72%) relative to the ASAP, and efficiency improvements are largely lost.
-This highlights that Stage 2—our surprisal-based pruning mechanism—is central to precisely identifying and preserving the most critical logical steps.
-In summary, the two stages operate in synergy: the coarse-grained anchor-guided pruning preliminarily removes redundant steps, while the fine-grained surprisal-based pruning further extracts the logical core of the reasoning trace.
-This two-stage process is key to ASAP’s strong performance, significantly outperforming any single-stage variant.
-Variants Acc $\uparrow$ Tok $\downarrow$ Lat $\downarrow$
-ASAP 36.19 6035 2.61
-w/o Coarse-grained Pruning 35.07 7735 4.60
-w/o Fine-grained Pruning 31.72 8061 4.83
-w/o Any Pruning 30.97 8289 4.83
-Table 2: Ablation study of different pruning strategies for ASAP on LiveCodeBench v4_v5.
-We report accuracy (Acc), average number of generated tokens (Tok), and average generation latency (Lat) measured in seconds.
-4.4 Impact of Token Budget
-To evaluate the performance scalability and resource sensitivity of our method, we analyze its behavior under varying inference-time token budgets (i.e., the maximum number of tokens the model is allowed to generate).
-We compare ASAP against the three strong baselines—SPIRIT, Original, and Zero-shot—across all benchmarks, and observe consistent trends.
-For clarity, we present results on LiveCodeBench v4_v5 under six budget settings ranging from 2K to 12K tokens.
-Results for other benchmarks and additional efficiency statistics are provided in Appendix C.
-As shown in Figure 3, our findings highlight three key advantages of ASAP: First, ASAP consistently outperforms all baselines across all budget settings on the benchmark.
-This demonstrates the robustness of our method under different resource constraints.
-Second, ASAP exhibits smooth performance scaling with respect to the token budget.
-Unlike baseline methods such as Zero-shot, whose accuracy can fluctuate as the budget increases (e.g., Zero-shot shows a decline from 8K to 10K), ASAP reliably improves with more available tokens.
-This predictable scaling behavior enables users to confidently configure budget limits to meet specific accuracy or latency targets.
-Third, ASAP achieves superior performance-efficiency trade-offs.
-For example, ASAP with just an 8K token budget achieves higher accuracy than SPIRIT and Original at a much larger 12K budget.
-These results further validate the practical utility of ASAP in real-world scenarios with constrained resources and varying performance requirements.
-Figure 3: Performance of ASAP on LiveCodeBench v4_v5 under different token budgets.
-4.5 Training Efficiency
-A central motivation for CoT compression is to reduce the computational burden during training and inference, specifically shortening the lengthy reasoning traces typically required by large reasoning models.
-Fine-tuning on compressed CoTs not only lowers training-time resource consumption but also helps the model internalize efficient reasoning patterns.
-To this end, we aim to achieve a high compression ratio across all methods while preserving the core logical steps.
-More details are provided in Appendix A.
-To quantify the training efficiency gains, we report two key metrics in Table 3: the average number of tokens per sample and the average training time measured in seconds per step.
-The results highlight the training efficiency advantage of ASAP.
-By generating the most compact yet logically rich CoTs, our approach significantly reduces training overhead.
-Compared to the uncompressed baseline (Original), our method reduces the number of training tokens by 75.6% and shortens training time by 60.7%.
-These savings substantially exceed those achieved by all other baselines.
-ASAP not only achieves a more favorable accuracy-efficiency trade-off at inference time, but also enables a more resource-efficient training process, making it a practical and cost-effective solution for real-world deployment.
-Methods Tokens Time
-Original 13023 80.11
-Selective Context 6722 (-48.4%) 63.41 (-20.9%)
-LLMLingua-2 6919 (-46.9%) 65.25 (-18.6%)
-TokenSkip 9813 (-24.6%) 77.27 (-3.6%)
-SPIRIT 6082 (-53.3%) 57.45 (-28.3%)
-ASAP 3178 (-75.6%) 31.48 (-60.7%)
-Table 3: Training efficiency comparison.
-We report the average number of tokens per sample and training time measured in seconds per step.
-Percentages indicate the reduction relative to the Original baseline.
-4.6 Generalization to Different Architectures
-To validate the generalizability of ASAP, we replicate our main experiments on the DeepSeek-R1-Distill-Llama-8B.
-Following the same experimental protocol, we compare ASAP against three strong baselines: Zero-shot, Original, and SPIRIT.
-We observe consistent trends across all benchmarks, so for brevity, we present representative results on two key benchmarks: LiveCodeBench v4_v5 and LeetCodeDataset in Table 4.
-The results across all benchmarks are reported in the Appendix D.
-The results on DeepSeek-R1-Distill-Llama-8B are highly consistent with our findings on the DeepSeek-R1-Distill-Qwen-7B, confirming that the effectiveness of ASAP generalizes beyond a specific model family.
-As shown in the Table 4, ASAP achieves the highest accuracy on both benchmarks, and the improvements in efficiency are even more pronounced.
-On LiveCodeBench, for instance, ASAP not only surpasses the accuracy of the Original baseline (32.84% vs.
-31.34%) but also generates 49.1% fewer tokens and reduces latency by over 3x (from 8.60s to 2.69s).
-Compared to the Zero-shot baseline, our method reduces token generation by 50.9% and latency by a remarkable 69.8%.
-This suggests that ASAP is particularly effective in identifying and distilling the core reasoning patterns, validating its robustness and broad applicability for improving reasoning efficiency across different model families.
-Methods LiveCodeBench LeetCodeDataset
-Acc $\uparrow$ Tok $\downarrow$ Lat $\downarrow$ Acc $\uparrow$ Tok $\downarrow$ Lat $\downarrow$
-Zero-shot 25.00 8508 8.90 27.19 8358 8.65
-Original 31.34 8202 8.60 26.32 8413 8.85
-SPIRIT 30.22 7913 8.45 26.75 8449 8.73
-ASAP 32.84 4175 2.69 27.63 3792 2.42
-Table 4: Experimental results of different methods with DeepSeek-R1-Distill-Llama-8B.
-We report accuracy (Acc), average number of generated tokens (Tok), and average generation latency (Lat) measured in seconds.
-The best results are highlighted in bold.
-5 Conclusion
-In this paper, we addressed the challenge of balancing performance and efficiency in Chain-of-Thought (CoT) reasoning, where longer reasoning traces improve accuracy but introduce significant redundancy and computational overhead.
-We proposed ASAP, a novel two-stage pruning framework that first generates an anchor for coarse-grained pruning, followed by a fine-grained refinement based on first-token surprisal.
-Extensive experiments demonstrate that models fine-tuned on CoTs pruned by ASAP achieve state-of-the-art results on multiple code generation benchmarks while substantially reducing both training and inference costs.
-Although our current work focuses on competitive programming tasks within the code generation domain, we believe the framework is general and broadly applicable.
-Appendix A Implementation Details
-Software and Hardware
+Appendix B Prompt Templates and Generated Examples
+We prompt the LLM to produce a direct thought: a concise, structured, step-by-step explanation that outlines how to derive the answer from the given question.
+A direct thought always consists of a small number of clear logical steps that directly contribute to the solution, and terminates with a single, explicit final answer.
+We use the prompt below to elicit such direct thoughts (the model is instructed to return only a detailed step-by-step solution containing only “Step-by-Step Solution" and “Final Answer").
+Prompt for Generating the Direct CoT $\mathcal{P}$ Given a question, please tell me how to get this answer step by step.
+Question: ``` {question} ``` Answer: ```python {answer} ``` Only return a detailed step-by-step solution (containing only “Step-by-Step Solution" and “Final Answer").
+The detailed step-by-step solution is:
+The following example illustrates a direct thought for a simple programming task: converting an integer into its binary representation.
+This example demonstrates the defining characteristics of direct CoTs, namely, concise step-by-step reasoning that directly derives the solution, without redundant exploration or verbose commentary.
+Example of a Direct CoT Step-by-Step Solution 1.
+Understand the Problem: The task is to convert a given positive integer n into its binary representation without leading zeros.
+The constraints are lenient ($n\leq 10^{6}$), so even a straightforward approach is efficient.
+2.
+Input Handling: Read the integer n from standard input.
+Since the input consists of a single integer, we can directly use input() and convert it to an integer.
+3.
+Binary Conversion: Python’s built-in bin() function converts an integer to its binary string representation (prefixed with 0b).
+To remove the prefix, slice the string from the third character onward ([2:]).
+4.
+Output the Result: Print the resulting binary string.
+This ensures no leading zeros are present.
+5.
+Efficiency: The approach is optimal because conversion and slicing both run in $O(\log n)$ time, and memory usage is minimal.
+Final Code: ```python n = int(input()) print(bin(n)[2:]) ```
+Given the direct thought as a reference, this stage instructs the model to prune redundant reasoning from the original chain of thought while preserving the essential logical structure.
+Specifically, the model is required to 1) remove unnecessary reasoning steps from the original CoT, 2) retain all key supporting content that aligns with the logic of direct CoT, and 3) strictly preserve the original wording and sentence order without introducing new information.
+This ensures that the compressed reasoning remains faithful to the original thought process while aligning with the concise, goal-oriented structure of the direct CoT.
+The following detailed prompt is used to elicit such coarse-grained pruning behavior.
+Prompt for Coarse-grained Pruning Compress the given thinking by referring to the provided solution.
+The goal is to remove irrelevant reasoning paths while retaining all content along the core reasoning path.
+Compression must be based on thinking, ensuring that the original wording and structure are preserved as much as possible.
+Follow these strict rules: 1.
+Use thinking as the foundation: Do not rewrite or replace its content with solution——only use solution to determine which parts are relevant.
+2.
+Remove unnecessary reasoning: Aggressively remove alternative paths that are not part of the core reasoning path.
+3.
+Retain key supporting content: Keep examples, reflections, and tests that help illustrate, verify, or analyze the core reasoning path.
+4.
+Preserve original words: Do not paraphrase, reorder, or change any words.
+5.
+Do not add new words: Do not introduce new concepts, symbols, or abbreviations.
+If you understand, compress the following thinking based on the given solution.
+Solution: ``` {solution} ``` Thinking: ``` {think} ``` The compressed thinking is:
+Appendix C Implementation Details
+Software and Hardware.
 For fine-tuning, we utilized the unsloth library [^1] https://pypi.org/project/unsloth/2025.5.6/ for its memory-efficient optimizations.
@@
 All experiments were conducted on NVIDIA H20 GPUs and Intel Xeon Platinum 8480+ CPUs.
-Fine-tuning Configuration
+Fine-tuning Configuration.
 We performed full-parameter fine-tuning for all models in our experiments.
@@
 Due to the high computational cost of full-parameter fine-tuning, the model is fine-tuned by a single run with a fixed random seed 42.
-Inference and Evaluation Protocol
+Inference and Evaluation Protocol.
 All inference benchmarks were run using the vLLM engine with dtype set to bfloat16 and gpu_memory_utilization set to 0.9.
 To ensure deterministic and reproducible outputs, we set the sampling temperature to 0.0 and set enable_prefix_caching to False.
-Baseline Details
+The default token budget for generation is adjusted based on the task difficulty.
+Specifically, it is 2K for GSM8K, 4K for MATH500, 6K for HumanEval+, and 10K for AIME24, AIME25, LiveCodeBench, and LeetCodeDataset.
+Results with other token budget settings are shown in Appendix F.
+Baseline Details.
 Following established practices, we used a consistent scoring model; as our primary model is DeepSeek-R1-Distill checkpoints, we employed DeepSeek-R1-Distill-Qwen-7B for all model-scoring tasks.
@@
 This variant retains the core idea of SPIRIT while improving scalability in our evaluation setting.
-Hyperparameters for Our Method
+Hyperparameters for Our Method.
 Our method involves several stages.
 For the LLM-guided Coarse-grained Pruning stage, we employed DeepSeek-V3 for economic reasons.
-When generating the Direct CoT, we used a deterministic setting (temperature=0.0, top_p=1.0), while for making the final pruning result, we increased exploration (temperature=2.0, top_p=1.0).
+When generating the direct thought $\mathcal{P}$, we used a deterministic setting (temperature=0.0, top_p=1.0), while for making the final pruning result, we increased exploration (temperature=1.0, top_p=1.0).
 For Pattern Matching, the similarity threshold $\tau$ was set to 0.6.
 Finally, during Surprisal-based Fine-grained Pruning, the maximum token budget was set to 4096 to ensure a deep level of compression.
-Appendix B Ablation Study of Different Pruning Strategies
+Appendix D Effect of Different Components.
 To validate the contribution and necessity of each component in our two-stage pruning framework, we conduct a detailed ablation study.
-Specifically, we evaluate the following three variants: ASAP w/o Coarse-grained Pruning, ASAP w/o Fine-grained Pruning and ASAP w/o Any Pruning.
-We present results on the HumanEval, HumanEval+, LiveCodeBench v1_v3, and LeetCodeDatsets benchmarks in Table 5, Table 6, Table 7, and Table 8.
-Variants Acc $\uparrow$ Tok $\downarrow$ Lat $\downarrow$
-ASAP 84.15 2464 0.98
-w/o Coarse-grained Pruning 82.32 2839 1.10
-w/o Fine-grained Pruning 71.34 2897 1.10
-w/o Any Pruning 78.05 2973 1.12
-Table 5: Ablation study of different pruning strategies for ASAP on HumanEval.
-We report accuracy (Acc), average number of generated tokens (Tok), and average generation latency (Lat) measured in seconds.
+Specifically, we evaluate the following three variants: ASAP w/o Coarse-grained Pruning, ASAP w/o Fine-grained Pruning, and ASAP w/o Any Pruning.
+We present results on the HumanEval+, LiveCodeBench v1_v3, and LeetCodeDatsets benchmarks in Table 6, Table 7, and Table 8.
 Variants Acc $\uparrow$ Tok $\downarrow$ Lat $\downarrow$
@@
 We report accuracy (Acc), average number of generated tokens (Tok), and average generation latency (Lat) measured in seconds.
-Appendix C Performance under Different Token Budgets
+Appendix E Generalization to Different Architectures
+To evaluate the generalizability of ASAP, we replicate our main experiments on the DeepSeek-R1-Distill-Llama-8B.
+Following the same experimental protocol, we compare ASAP against three baselines: Zero-shot, Original, and SPIRIT.
+The results of the code generation task on the HumanEval+, LiveCodeBench v1_v3, LiveCodeBench v4_v5, and LeetCodeDataset benchmarks are shown in Table 9.
+The results of the mathematical reasoning task on the GSM8K, MATH500, AIME24, and AIME25 benchmarks are shown in Table 10.
+Methods HE+ LCBv1_v3 LCBv4_v5 LCD
+Acc $\uparrow$ Tok $\downarrow$ Lat $\downarrow$ Acc $\uparrow$ Tok $\downarrow$ Lat $\downarrow$ Acc $\uparrow$ Tok $\downarrow$ Lat $\downarrow$ Acc $\uparrow$ Tok $\downarrow$ Lat $\downarrow$
+Zero-shot 64.02 3334 1.86 44.12 7162 6.92 25.00 8508 8.90 27.19 8358 8.65
+Original 76.22 2978 1.63 52.61 6614 6.16 31.34 8202 8.60 26.32 8413 8.85
+SPIRIT 72.56 3159 1.74 52.61 6280 5.84 30.22 7913 8.45 26.75 8449 8.73
+ASAP 76.83 2494 1.30 48.86 3605 2.18 32.84 4175 2.69 27.63 3792 2.42
+Table 9: Experimental results of different methods on code generation benchmarks with DeepSeek-R1-Distill-Llama-8B.
+We report accuracy (Acc), average number of generated tokens (Tok), and average generation latency (Lat) measured in seconds.
+Methods GSM8K MATH500 AIME24 AIME25
+Acc $\uparrow$ Tok $\downarrow$ Lat $\downarrow$ Acc $\uparrow$ Tok $\downarrow$ Lat $\downarrow$ Acc $\uparrow$ Tok $\downarrow$ Lat $\downarrow$ Acc $\uparrow$ Tok $\downarrow$ Lat $\downarrow$
+Zero-shot 79.15 1262 0.36 57.20 2612 1.08 33.33 8445 10.42 26.67 8597 10.54
+Original 84.91 1310 0.37 63.00 2534 1.01 36.67 8550 10.04 30.00 8268 10.05
+SPIRIT 85.67 1256 0.35 62.60 2533 1.01 36.67 8788 10.04 36.67 8094 9.57
+ASAP 87.34 768 0.20 66.00 1734 0.65 36.67 5314 6.97 33.33 5348 7.05
+Table 10: Experimental results of different methods on mathematical reasoning benchmarks with DeepSeek-R1-Distill-Llama-8B.
+We report accuracy (Acc), average number of generated tokens (Tok), and average generation latency (Lat) measured in seconds.
+Appendix F Performance under Different Token Budgets
 To evaluate the performance scalability and resource sensitivity of our method, we analyze its behavior under varying inference-time token budgets (i.e., the maximum number of tokens the model is allowed to generate).
-We compare ASAP with three strong baselines—SPIRIT, Original, and Zero-shot—on HumanEval, LiveCodeBench, and LeetcodeDataset.
-For HumanEval, we evaluate the performance under four budget settings, ranging from 1K to 6K tokens.
-For LiveCodeBench and LeetCodeDataset, we evaluate the performance under six budget settings, ranging from 2K to 12K tokens.
-Results are shown in Table 9, Table 10 and Table 11.
+We compare ASAP with three strong baselines—SPIRIT, Original, and Zero-shot—on HumanEval+, LiveCodeBench v1_v3, LiveCodeBench v4_v5, LeetcodeDataset, GSM8K, MATH500, AIME24, and AIME25.
+For simpler benchmarks (including HumanEval+, GSM8K, and MATH500), we evaluate the performance under four budget settings, ranging from 1K to 6K tokens.
+For more complex benchmarks (including LiveCodeBench v1_v3, LiveCodeBench v4_v5, LeetcodeDataset, AIME24, and AIME25), we evaluate the performance under six budget settings, ranging from 2K to 12K tokens.
+Results are shown in Table 11, Table 12, Table 13, Table 14, Table 15, Table 16, Table 17, and Table 18.
 Budget Zero-shot Original SPIRIT ASAP
-Acc Tok Lat Acc Tok Lat Acc Tok Lat Acc Tok Lat
-1K 10.37 1007 0.28 15.85 983 0.28 11.59 995 0.28 25.00 946 0.27
-2K 45.12 1813 0.53 44.51 1702 0.49 49.39 1690 0.49 56.71 1502 0.44
-4K 71.95 2561 0.85 67.68 2511 0.82 76.83 2401 0.80 75.00 2116 0.72
-6K 73.78 3051 1.16 78.05 2973 1.12 83.54 2764 1.07 84.15 2464 0.98
-(a) Results on HumanEval
-Budget Zero-Shot Original SPIRIT ASAP
-Acc Tok Lat Acc Tok Lat Acc Tok Lat Acc Tok Lat
+Acc $\uparrow$ Tok $\downarrow$ Lat $\downarrow$ Acc $\uparrow$ Tok $\downarrow$ Lat $\downarrow$ Acc $\uparrow$ Tok $\downarrow$ Lat $\downarrow$ Acc $\uparrow$ Tok $\downarrow$ Lat $\downarrow$
 1K 9.76 1007 0.28 14.63 983 0.28 10.98 995 0.28 23.78 946 0.27
@@
 6K 68.29 3051 1.16 75.61 2973 1.12 75.61 2764 1.07 78.66 2464 0.98
-(b) Results on HumanEval+
-Table 9: Results of different methods under different budgets on HumanEval.
+Table 11: Results of different methods under different budgets on HumanEval+.
 We report accuracy (Acc), average number of generated tokens (Tok), and average generation latency (Lat) measured in seconds.
 .
-Budget Zero-shot Original SPIRIT ASAP Acc Tok Lat Acc Tok Lat Acc Tok Lat Acc Tok Lat 2K 16.50 1966 0.52 17.16 1920 0.51 18.95 1908 0.51 21.57 1833 0.49 4K 32.68 3499 1.06 30.72 3432 1.05 34.80 3370 1.03 34.97 3244 1.00 6K 39.05 4806 1.70 42.65 4673 1.67 43.14 4605 1.64 46.24 4358 1.54 8K 44.28 5903 2.46 47.71 5723 2.43 51.80 5515 2.27 52.61 4919 1.90 10K 42.16 7088 3.59 52.12 6611 3.15 50.82 6524 3.09 54.74 5177 2.09 12K 43.95 7988 5.10 54.41 7473 4.22 51.63 7362 4.09 55.56 5322 2.27
-(a) Results on LiveCodeBench v1_v3
+Budget Zero-shot Original SPIRIT ASAP Acc $\uparrow$ Tok $\downarrow$ Lat $\downarrow$ Acc $\uparrow$ Tok $\downarrow$ Lat $\downarrow$ Acc $\uparrow$ Tok $\downarrow$ Lat $\downarrow$ Acc $\uparrow$ Tok $\downarrow$ Lat $\downarrow$ 2K 16.50 1966 0.52 17.16 1920 0.51 18.95 1908 0.51 21.57 1833 0.49 4K 32.68 3499 1.06 30.72 3432 1.05 34.80 3370 1.03 34.97 3244 1.00 6K 39.05 4806 1.70 42.65 4673 1.67 43.14 4605 1.64 46.24 4358 1.54 8K 44.28 5903 2.46 47.71 5723 2.43 51.80 5515 2.27 52.61 4919 1.90 10K 42.16 7088 3.59 52.12 6611 3.15 50.82 6524 3.09 54.74 5177 2.09 12K 43.95 7988 5.10 54.41 7473 4.22 51.63 7362 4.09 55.56 5322 2.27
+Table 12: Results of different methods under different budgets on LiveCodeBench v1_v3.
+We report accuracy (Acc), average number of generated tokens (Tok), and average generation latency (Lat) measured in seconds.
 Budget Zero-shot Original SPIRIT ASAP
-Acc Tok Lat Acc Tok Lat Acc Tok Lat Acc Tok Lat
+Acc $\uparrow$ Tok $\downarrow$ Lat $\downarrow$ Acc $\uparrow$ Tok $\downarrow$ Lat $\downarrow$ Acc $\uparrow$ Tok $\downarrow$ Lat $\downarrow$ Acc $\uparrow$ Tok $\downarrow$ Lat $\downarrow$
 2K 6.72 2021 0.59 6.34 1999 0.57 8.21 1993 0.56 13.43 1930 0.54
@@
 12K 25.75 9706 7.44 32.46 9567 7.10 34.33 8987 6.73 36.57 6128 2.76
-(b) Results on LiveCodeBench v4_v5
-Table 10: Results of different methods under different budgets on LiveCodeBench.
+Table 13: Results of different methods under different budgets on LiveCodeBench v4_v5.
 We report accuracy (Acc), average number of generated tokens (Tok), and average generation latency (Lat) measured in seconds.
 Budget Zero-shot Original SPIRIT ASAP
-Acc Tok Lat Acc Tok Lat Acc Tok Lat Acc Tok Lat
+Acc $\uparrow$ Tok $\downarrow$ Lat $\downarrow$ Acc $\uparrow$ Tok $\downarrow$ Lat $\downarrow$ Acc $\uparrow$ Tok $\downarrow$ Lat $\downarrow$ Acc $\uparrow$ Tok $\downarrow$ Lat $\downarrow$
 2K 7.02 2028 0.53 6.14 2020 0.53 7.02 2001 0.53 10.09 1965 0.53
@@
 We report accuracy (Acc), average number of generated tokens (Tok), and average generation latency (Lat) measured in seconds.
-Appendix D Generalization to Different Architectures
-To evaluate the generalizability of ASAP, we replicate our main experiments on the DeepSeek-R1-Distill-Llama-8B.
-Following the same experimental protocol, we compare ASAP against three baselines: Zero-shot, Original, and SPIRIT.
-Results on the HumanEval, HumanEval+, LiveCodeBench v1_v3, LiveCodeBench v4_v5, and LeetCodeDataset benchmarks are shown in Table 12.
-Methods HumanEval LiveCodeBench LeetCodeDataset
-HE HE+ v1_v3 v4_v5 All
-Acc $\uparrow$ Tok $\downarrow$ Lat $\downarrow$ Acc $\uparrow$ Tok $\downarrow$ Lat $\downarrow$ Acc $\uparrow$ Tok $\downarrow$ Lat $\downarrow$ Acc $\uparrow$ Tok $\downarrow$ Lat $\downarrow$ Acc $\uparrow$ Tok $\downarrow$ Lat $\downarrow$
-Zero-shot 68.29 3334 1.86 64.02 3334 1.86 44.12 7162 6.92 25.00 8508 8.90 27.19 8358 8.65
-Original 82.32 2978 1.63 76.22 2978 1.63 52.61 6614 6.16 31.34 8202 8.60 26.32 8413 8.85
-SPIRIT 76.22 3159 1.74 72.56 3159 1.74 52.61 6280 5.84 30.22 7913 8.45 26.75 8449 8.73
-ASAP 80.49 2494 1.30 76.83 2494 1.30 48.86 3605 2.18 32.84 4175 2.69 27.63 3792 2.42
-Table 12: Experimental results of different methods with DeepSeek-R1-Distill-Llama-8B.
-We report accuracy (Acc), average number of generated tokens (Tok), and average generation latency (Lat) measured in seconds.
-Appendix E Prompts Used in Our Method
-Prompt for Direct CoT Generation Given a question, please tell me how to get this answer step by step.
-Question: ``` {question} ``` Answer: ```python {answer} ``` Only return a detailed (with code) step-by-step solution (containing only “Step-by-Step Solution” and “Final Code”).
-The detailed step-by-step solution is:
-Prompt for Coarse-grained Pruning Compress the given thinking by referring to the provided solution.
-The goal is to remove irrelevant reasoning paths while retaining all content along the core reasoning path.
-Compression must be based on thinking, ensuring that the original wording and structure are preserved as much as possible.
-Follow these strict rules: 1.
-Use thinking as the foundation: Do not rewrite or replace its content with solution——only use solution to determine which parts are relevant.
-2.
-Remove unnecessary reasoning: Aggressively remove alternative paths that are not part of the core reasoning path.
-3.
-Retain key supporting content: Keep examples, reflections, and tests that help illustrate, verify, or analyze the core reasoning path.
-4.
-Preserve original words: Do not paraphrase, reorder, or change any words.
-5.
-Do not add new words: Do not introduce new concepts, symbols, or abbreviations.
-If you understand, compress the following thinking based on the given solution.
-Solution: ``` {solution} ``` Thinking: ``` {think} ``` The compressed thinking is:
+Budget Zero-shot Original SPIRIT ASAP
+Acc $\uparrow$ Tok $\downarrow$ Lat $\downarrow$ Acc $\uparrow$ Tok $\downarrow$ Lat $\downarrow$ Acc $\uparrow$ Tok $\downarrow$ Lat $\downarrow$ Acc $\uparrow$ Tok $\downarrow$ Lat $\downarrow$
+1K 48.75 963 0.19 59.29 942 0.19 54.66 925 0.18 83.93 693 0.14
+2K 83.55 1301 0.27 86.35 1250 0.26 88.55 1118 0.23 90.75 753 0.16
+4K 88.65 1553 0.37 90.37 1432 0.34 90.52 1227 0.28 91.28 778 0.18
+6K 89.23 1714 0.46 91.05 1513 0.39 91.28 1297 0.33 91.81 790 0.20
+Table 15: Results of different methods under different budgets on GSM8K. We report accuracy (Acc), average number of generated tokens (Tok), and average generation latency (Lat) measured in seconds.
+Budget Zero-shot Original SPIRIT ASAP
+Acc $\uparrow$ Tok $\downarrow$ Lat $\downarrow$ Acc $\uparrow$ Tok $\downarrow$ Lat $\downarrow$ Acc $\uparrow$ Tok $\downarrow$ Lat $\downarrow$ Acc $\uparrow$ Tok $\downarrow$ Lat $\downarrow$
+1K 19.00 1020 0.19 28.00 1017 0.19 18.40 1012 0.19 36.40 935 0.19
+2K 42.20 1804 0.39 52.00 1767 0.39 54.40 1592 0.36 59.80 1347 0.31
+4K 60.40 2629 0.70 63.80 2511 0.66 64.20 2144 0.57 70.80 1649 0.43
+6K 66.60 3100 0.94 70.60 2843 0.84 69.60 2460 0.74 71.00 1758 0.52
+Table 16: Results of different methods under different budgets on MATH500. We report accuracy (Acc), average number of generated tokens (Tok), and average generation latency (Lat) measured in seconds.
+.
+Budget Zero-shot Original SPIRIT ASAP Acc $\uparrow$ Tok $\downarrow$ Lat $\downarrow$ Acc $\uparrow$ Tok $\downarrow$ Lat $\downarrow$ Acc $\uparrow$ Tok $\downarrow$ Lat $\downarrow$ Acc $\uparrow$ Tok $\downarrow$ Lat $\downarrow$ 2K 0.00 2048 0.93 16.67 2048 0.93 6.67 2048 0.93 10.00 1978 0.93 4K 20.00 4003 2.12 20.00 3984 2.11 23.33 3877 2.10 36.67 3415 1.99 6K 33.33 5682 3.58 36.67 5695 3.55 40.00 5216 3.32 36.67 4410 3.03 8K 30.00 7073 5.30 40.00 7093 5.14 46.67 6243 4.47 40.00 5159 4.10 10K 36.67 8352 6.76 46.67 8034 6.43 46.67 7198 5.78 46.67 5552 5.04 12K 40.00 9318 7.84 46.67 8990 7.75 46.67 8363 7.21 46.67 5767 5.88
+Table 17: Results of different methods under different budgets on AIME24. We report accuracy (Acc), average number of generated tokens (Tok), and average generation latency (Lat) measured in seconds.
+.
+Budget Zero-shot Original SPIRIT ASAP Acc $\uparrow$ Tok $\downarrow$ Lat $\downarrow$ Acc $\uparrow$ Tok $\downarrow$ Lat $\downarrow$ Acc $\uparrow$ Tok $\downarrow$ Lat $\downarrow$ Acc $\uparrow$ Tok $\downarrow$ Lat $\downarrow$ 2K 6.67 2046 0.95 13.33 2048 0.94 3.33 2044 0.94 10.00 2020 0.95 4K 20.00 3851 2.17 26.67 3834 2.16 16.67 3792 2.15 20.00 3511 2.15 6K 30.00 5369 3.62 33.33 5452 3.63 36.67 5360 3.63 30.00 4484 3.14 8K 36.67 6848 5.30 36.67 6798 5.32 36.67 6611 5.16 33.33 5002 4.10 10K 40.00 8145 6.67 43.33 8026 6.75 43.33 7817 6.57 36.67 5434 5.10 12K 36.67 9442 8.25 40.00 9461 8.32 46.67 8598 7.78 36.67 5720 6.06
+Table 18: Results of different methods under different budgets on AIME25. We report accuracy (Acc), average number of generated tokens (Tok), and average generation latency (Lat) measured in seconds.
+Appendix G Training Efficiency
+To quantify the training efficiency gains, we present results of the CodeForces-CoTs dataset in Table 5 and results of the OpenR1-Math dataset in Table 19.
+We report two key metrics: the average number of tokens per sample and the average training time measured in seconds per step.
+Methods Tokens Time
+Original 5807 47.82
+Selective Context 3149 (-45.8%) 25.85 (-45.9%)
+LLMLingua-2 3478 (-40.1%) 28.75 (-39.9%)
+TokenSkip 4728 (-18.6%) 39.20 (-18.0%)
+SPIRIT 2858 (-50.8%) 23.67 (-50.5%)
+ASAP 1834 (-68.4%) 15.36 (-67.9%)
+Table 19: Training efficiency comparison on OpenR1-Math dataset.
+We report the average number of tokens per sample and training time measured in seconds per step.
+Percentages indicate the reduction relative to the Original baseline.